### v2

# Tilting representations of finite groups of Lie type

Abstract. Let $\mathbf{G}$ be a connected reductive group over a finite field $\mathbb{F}_{q}$ of characteristic $p>0$. In this paper, we study a category which we call Deligne–Lusztig category $\mathcal{O}$ and whose definition is similar to category $\mathcal{O}$. We use this to construct a collection of representations of $\mathbf{G}(\mathbb{F}_{q})$ which we call the tilting representations. They form a generating collection of integral projective representations of $\mathbf{G}(\mathbb{F}_{q})$. Finally we compute the character of these representations and relate their expression to previous calculations of Lusztig and we then use this to establish a conjecture of Dudas–Malle of 2014.
1. Introduction
Let $p>0$ be a prime and $q=p^{r}$ a power of $p$ and let $\mathbb{F}_{q}\subset\overline{\mathbb{F}}_{q}=k$ be the finite field with $q$ elements and its algebraic closure. Let $\mathbf{G}$ be a connected reductive group over $k$ equipped with a Frobenius endomorphism $\mathrm{F}:\mathbf{G}\to\mathbf{G}$, i.e. a purely inseparable isogeny such that a power of $\mathrm{F}$ is the Frobenius coming from an $\mathbb{F}_{q^{s}}$-structure. We fix $\mathbf{B}=\mathbf{T}\mathbf{U}$ a Borel pair that is stable under $\mathrm{F}$ and we denote by $\mathbf{W}$ the corresponding Weyl group. This paper has two main goals :
  1. $(i)$.
    Introduce and study a category which we call ‘Deligne–Lusztig category $\mathcal{O}$’ whose definition mimics the geometric realization of classical category $\mathcal{O}$ as sheaves on $\mathbf{U}\backslash\mathbf{G}/\mathbf{B}$.
  2. $(ii)$.
    Introduce a class of representations of the finite group of Lie type ${\mathbf{G}}^{\mathrm{F}}$, which we call tilting representations and study their properties.
1.1. Deligne–Lusztig category $\mathcal{O}$
1.1.1. Category $\mathcal{O}$
Category $\mathcal{O}$ is a central object in representation theory, it links, through the Kazhdan–Lusztig conjectures [KL79], [KL80], the theory of representations of semisimple Lie algebras with the geometry of the flag varieties and provides a categorification of the Hecke algebra associated to $\mathbf{W}$, we refer to [Ach21] for a historical account. It is now standard to define
$\mathcal{O}=\mathrm{D}^{b}_{c,\mathbf{U}}(\mathbf{G}/\mathbf{B},\overline{\mathbb{Q}}_{\ell})$
as the category of $\ell$-adic sheaves, where $\ell\neq p$ is a prime number, on the flag variety of $\mathbf{G}$ that are locally constant along $\mathbf{U}$-orbits. Its perverse heart is an abelian category which has the structure of a highest weight category [BGS96] and therefore has very good homological properties. While it is known that this category provides a categorification of the Hecke algebra, it is not equipped with a monoidal structure. to reveal the algebra structure of the Hecke algebra, one must pass to Hecke categories. They come in two flavors.
  1. $(i)$.
    The equivariant Hecke category: $\mathcal{H}^{\mathrm{eq}}$. We define it, for this introduction, as the category of sheaves
    $\mathrm{D}^{b}_{c,\mathbf{B}}(\mathbf{G}/\mathbf{B},\overline{\mathbb{Q}}_{\ell})=\mathrm{D}^{b}_{c}(\mathbf{B}\backslash\mathbf{G}/\mathbf{B},\overline{\mathbb{Q}}_{\ell})$
    the category of $\mathbf{B}$-equivariant sheaves on $\mathbf{G}/\mathbf{B}$. This is a monoidal category that acts on category $\mathcal{O}$ which categorifies the Hecke algebra.
  2. $(ii)$.
    The monodromic Hecke category : $\mathcal{H}^{\mathrm{mon}}$, initially constructed by [BY13]. It is a certain category of sheaves on $\mathbf{U}\backslash\mathbf{G}/\mathbf{U}.$ It follows from loc. cit. that this category also provides a categorification of the Hecke algebra and acts on category $\mathcal{O}$.
Both of these actions categorify the left and right module structures of the Hecke algebra on itself. We equip $\mathbf{G}$ with its Bruhat stratification $\mathbf{G}=\sqcup_{w\in\mathbf{W}}\mathbf{B}w\mathbf{B}$, this induces a stratification of $\mathbf{G}/\mathbf{B}$ and all the quotients mentioned before. We let $j_{w}:\mathbf{B}w\mathbf{B}/\mathbf{B}\subset\mathbf{G}/\mathbf{B}$ be the inclusion. Category $\mathcal{O}$ possesses four collections of perverse sheaves that are of interest to representation theory :
  1. $(i)$.
    the standard sheaves $\Delta_{w}^{\mathcal{O}}=j_{w,!}\overline{\mathbb{Q}}_{\ell}[\ell(w)]$,
  2. $(ii)$.
    the costandard sheaves $\nabla_{w}^{\mathcal{O}}=j_{w,*}\overline{\mathbb{Q}}_{\ell}[\ell(w)]$,
  3. $(iii)$.
    the simple sheaves $\mathrm{IC}_{w}^{\mathcal{O}}=j_{w,!*}\overline{\mathbb{Q}}_{\ell}[\ell(w)]$,
  4. $(iv)$.
    the indecomposable tilting sheaves $T_{w}^{\mathcal{O}}$, where $\ell(w)$ is the length of the element $w\in\mathbf{W}$.
The tilting objects are defined as follows.
Definition 1.1.1.
Let $A\in\mathrm{Perv}_{\mathbf{U}}(\mathbf{G}/\mathbf{B},\overline{\mathbb{Q}}_{\ell})$. Then
  1. $(i)$.
    a $\Delta$-filtration on $A$ is a filtration whose graded pieces are standard sheaves,
  2. $(ii)$.
    a $\nabla$-filtration on $A$ is a filtration whose graded pieces are costandard sheaves,
  3. $(iii)$.
    the object $A$ is tilting if it has both a $\Delta$ and a $\nabla$-filtration.
In the context of category $\mathcal{O}$, it is known that there exists a unique indecomposable tilting sheaf $T_{w}^{\mathcal{O}}$ such that $T_{w}^{\mathcal{O}}$ contains $\Delta_{w}^{\mathcal{O}}$ in some (equivalently any) $\Delta$-filtration with multiplicity one and $T_{w}^{\mathcal{O}}$ is supported on the closure of $\mathbf{B}w\mathbf{B}/\mathbf{B}$.
Hecke categories only have some of these collections, namely
  1. $(i)$.
    the equivariant Hecke category $\mathcal{H}^{\mathrm{eq}}$ has standard, costandard and simple objects but usually no tilting objects,
  2. $(ii)$.
    the monodromic Hecke category $\mathcal{H}^{\mathrm{mon}}$ has standard, costandard and tilting objects but usually not a good theory of simple object.
So far, we have restricted ourselves to the case of $\overline{\mathbb{Q}}_{\ell}$-sheaves but much of what has been discussed so far holds with $\overline{\mathbb{F}}_{\ell}$-sheaves instead, see for instance [AR16], [RSW14].
1.1.2. Deligne–Lusztig catgory $\mathcal{O}$
In our context we also have the Frobenius endomorphism $\mathrm{F}:\mathbf{G}\to\mathbf{G}$. We are interested in the representation theory of the finite group ${\mathbf{G}}^{\mathrm{F}}$. Consider the action of $\mathbf{G}$ on itself by $g.x=gx\mathrm{F}(g^{-1})$, we denote this action by $\mathrm{Ad}_{\mathrm{F}}$. Since $\mathbf{G}$ is connected Lang’s theorem provides an isomorphism of algebraic stacks
$\frac{\mathbf{G}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{G}}=\mathrm{pt}/{\mathbf{G}}^{\mathrm{F}}$
and by étale descent there is a natural equivalence of categories $\mathrm{D}(\mathrm{Rep}_{\Lambda}{\mathbf{G}}^{\mathrm{F}})=\mathrm{D}(\mathrm{pt}/{\mathbf{G}}^{\mathrm{F}},\Lambda)$, where $\Lambda$ is a coefficient ring for which the category of étale sheaves on $\mathrm{pt}/{\mathbf{G}}^{\mathrm{F}}$ makes sense, see the notation section 1.3. We introduce the stack
$\frac{\mathbf{U}\backslash\mathbf{G}/\mathbf{U}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{T}}$
which we call the horocycle stack. We had previously introduced this stack in [Ete24]. We define
$\mathcal{O}^{\mathrm{DL}}_{\Lambda}=\mathrm{D}(\frac{\mathbf{U}\backslash\mathbf{G}/\mathbf{U}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{T}},\Lambda)$
the category of $\Lambda$-sheaves on it where $\Lambda\in\{\overline{\mathbb{F}}_{\ell},\overline{\mathbb{Q}}_{\ell},\overline{\mathbb{Z}}_{\ell}\}$. We call this category ‘Deligne–Lusztig category $\mathcal{O}$’. We consider it as a variant of the classical category $\mathcal{O}$. To put some emphasis on the comparison, category $\mathcal{O}$ is classically realized as the category of sheaves on
$(\mathbf{U}\backslash\mathbf{G}/\mathbf{U})/\mathbf{T}$
where the torus $\mathbf{T}$ acts by right translations.
We equip this stack with its Bruhat stratification and, for $w\in\mathbf{W}$, the corresponding stratum is then isomorphic to
$\frac{\mathbf{U}\backslash\mathbf{B}w\mathbf{B}/\mathbf{U}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{T}}=\mathrm{pt}/({\mathbf{T}}^{w\mathrm{F}}\ltimes\mathbf{U}\cap{{}^{w}}\mathbf{U}).$
We refer to section 2 for a discussion. There is a natural equivalence $\mathrm{D}(\mathrm{pt}/({\mathbf{T}}^{w\mathrm{F}}\ltimes\mathbf{U}\cap{{}^{w}}\mathbf{U}),\Lambda)=\mathrm{D}(\mathrm{Rep}_{\Lambda}{\mathbf{T}}^{w\mathrm{F}})$ which we normalize to be $t$-exact. It follows that $\mathcal{O}^{\mathrm{DL}}$ is glued from the categories of representations of the finite groups ${\mathbf{T}}^{w\mathrm{F}}$. Given a character $\chi:{\mathbf{T}}^{w\mathrm{F}}\to\Lambda^{\times}$ of ${\mathbf{T}}^{w\mathrm{F}}$, let $E_{\chi}$ be the projective cover of $\chi$ in the abelian category $\mathrm{Rep}_{\Lambda}{\mathbf{T}}^{w\mathrm{F}}$ of representations of ${\mathbf{T}}^{w\mathrm{F}}$ on $\Lambda$-modules. We set
  1. $(i)$.
    $\Delta_{w,\chi}=j_{w,!}E_{\chi}[\ell(w)]$, the standard objects
  2. $(ii)$.
    $\nabla_{w,\chi}=j_{w,*}E_{\chi}[\ell(w)]$, the costandard objects.
We denote by $\mathcal{O}^{\mathrm{DL},\heartsuit}$ the perverse heart of $\mathcal{O}^{\mathrm{DL}}$.
We then proceed to study this category and we show that this category has much of the structure of a highest weight category. Let us summarize the main properties of this category, see Theorem 2.3.2 and Lemma 2.3.5.
Theorem 1.1.2.
  1. $(i)$.
    The category $\mathcal{O}^{\mathrm{DL}}$ is compactly generated, and the standard and costandard sheaves are compact.
  2. $(ii)$.
    For all pairs $(w,\chi)$, where $w\in\mathbf{W}$ and $\chi$ is a character of ${\mathbf{T}}^{w\mathrm{F}}$, there exists a unique (up to isomorphism) indecomposable tilting sheaf $T_{w,\chi}$ supported on the closure of $\mathbf{B}w\mathbf{B}$ containing $\Delta_{w,\chi}$ with multiplicity $1$ in a $\Delta$-filtration, moreover any indecomposable tilting sheaf is of this form.
  3. $(iii)$.
    A sheaf $T\in\mathcal{O}^{\mathrm{DL},\heartsuit}$ is tilting if and only if for all $w$, the objects $j_{w}^{*}T$ and $j_{w}^{!}T$ are perverse and projective.
  4. $(iv)$.
    The subcategory $\mathcal{O}^{\mathrm{DL}}_{\mathrm{tilt}}$ of tilting sheaves is an additive, Karoubian category satisfying the Krull–Schmidt property and generates $\mathcal{O}^{\mathrm{DL}}$.
We also discuss the following relations
  1. $(i)$.
    We compare $\mathcal{O}^{\mathrm{DL}}$ with the Hecke categories $\mathcal{H}^{\mathrm{eq}}$ and $\mathcal{H}^{\mathrm{mon}}$ and their twisted variants of [LY20] and [Gou22] in section $2.$
  2. $(ii)$.
    We define a mixed version $\mathcal{O}^{\mathrm{DL},\mathrm{mix}}_{\overline{\mathbb{Q}}_{\ell}}$ of $\mathcal{O}^{\mathrm{DL}}_{\overline{\mathbb{Q}}_{\ell}}$ and we show categorification results in the same vein as the categorification results of $\mathcal{O}$, see section 2.5.
  3. $(iii)$.
    We compare the decomposition numbers of tilting sheaves of $\mathcal{O}^{\mathrm{DL}}$ with the ones of $\mathcal{O}$ in section 2.6.
It follows from this discussion that the structure of $\mathcal{O}^{\mathrm{DL}}$ is essentially controlled by Kazhdan–Lusztig type combinatorics.
1.2. Tilting representations
1.2.1. Deligne–Lusztig theory
Classically, the representation theory of ${\mathbf{G}}^{\mathrm{F}}$ is studied using Deligne–Lusztig theory [DL76]. One first defines two collections of varieties
  1. $(i)$.
    given $w\in\mathbf{W}$, we define
    $X(w)=\{g\mathbf{B},g^{-1}\mathrm{F}(g)\in\mathbf{B}w\mathbf{B}\}\subset\mathbf{G}/\mathbf{B}$
  2. $(ii)$.
    and given a lift $\dot{w}\in\mathbf{N}(\mathbf{T})$ of the element $w\in\mathbf{W}$, we define
    $Y(\dot{w})=\{g\mathbf{U},g^{-1}\mathrm{F}(g)\in\mathbf{U}\dot{w}\mathbf{U}\}\subset\mathbf{G}/\mathbf{U}.$
These varieties are smooth of dimension $\ell(w)$ and are equipped with an action of the finite group ${\mathbf{G}}^{\mathrm{F}}$. Taking their cohomology yields representations of ${\mathbf{G}}^{\mathrm{F}}$ and one usually denotes by
$R_{w}^{1}=\sum_{i}(-1)^{i}[H^{i}_{c}(X(w),\overline{\mathbb{Q}}_{\ell})]$
the character of $\mathrm{R\Gamma}_{c}(X(w),\overline{\mathbb{Q}}_{\ell})$. More generally, since the variety $Y(\dot{w})$ also carries an action of ${\mathbf{T}}^{w\mathrm{F}}$, we also define
$R_{w}^{\theta}=\sum_{i}(-1)^{i}[H^{i}_{c}(Y(\dot{w}),\overline{\mathbb{Q}}_{\ell})]_{\theta}$
by taking $\theta$ isotypic parts, where $\theta$ is a character of ${\mathbf{T}}^{w\mathrm{F}}$. A very important set of class functions on ${\mathbf{G}}^{\mathrm{F}}$ are the uniform class functions. These are the class functions of ${\mathbf{G}}^{\mathrm{F}}$ that are linear combinations of the characters $R_{w}^{\theta}$. An interesting example is provided by the character
$\mathrm{IH}(\overline{X(w)},\overline{\mathbb{Q}}_{\ell})]=\sum_{i}(-1)^{i}[\mathrm{IH}^{i}(\overline{X(w)},\overline{\mathbb{Q}}_{\ell})]$
of the intersection cohomology of the closure of the variety $X(w)$ in $\mathbf{G}/\mathbf{B}$ (which is usually a singular variety). There is an explicit relation between this character and the character $R_{w}^{1}$ expressed in terms of Kazhdan–Lusztig polynomials done in [Lus84], this relation is a reflection of some geometric relation between Deligne–Lusztig theory and the Hecke category.
1.2.2. Tilting representations
There is a natural functor
$\mathrm{ch}=\alpha_{!}\beta^{*}[\dim\mathbf{U}]:\mathcal{O}^{\mathrm{DL}}\to\mathrm{D}(\mathrm{pt}/{\mathbf{G}}^{\mathrm{F}},\Lambda)$
where the maps $\alpha$ and $\beta$ are given by the following correspondence
$\frac{\mathbf{G}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{G}}\xleftarrow{\alpha}\frac{\mathbf{G}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{B}}\xrightarrow{\beta}\frac{\mathbf{U}\backslash\mathbf{G}/\mathbf{U}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{T}}.$
We refer to section 3. This correspondence is a variant of the correspondence used by Lusztig to define character sheaves [Lus85] and more recently in [Lus15]. A variant with $\mathbf{B}\backslash\mathbf{G}/\mathbf{B}$ instead of $\frac{\mathbf{U}\backslash\mathbf{G}/\mathbf{U}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{T}}$ was studied in [BDR20].
Theorem 1.2.1 ([Ete24]).
There are isomorphisms up to shifts
$\mathrm{ch}(\Delta_{w,\chi})=e_{\chi}\mathrm{R\Gamma}_{c}(Y(\dot{w}),\Lambda)[\ell(w)]$
and
$\mathrm{ch}(\nabla_{w,\chi})=e_{\chi}\mathrm{R\Gamma}(Y(\dot{w}),\Lambda)[\ell(w)]$
where $e_{\chi}$ is the projector for the action of ${\mathbf{T}}^{w\mathrm{F}}$ on the block containing $\chi$.
Remark 1.2.2.
It follows from this theorem that $\mathcal{O}^{\mathrm{DL}}$ is a very rich source of uniform functions. There is however a lot more information contained in this category than in the space of uniform class functions.
Definition 1.2.3.
A tilting representation of ${\mathbf{G}}^{\mathrm{F}}$ is a complex of representations of the form
$\mathrm{ch}(T)\in\mathrm{D}(\mathrm{pt}/{\mathbf{G}}^{\mathrm{F}},\Lambda)$
where $T\in\mathcal{O}^{\mathrm{DL}}$ is a perverse tilting object.
The next theorem has also been discovered independently by Zhu [Zhu25].
Theorem 1.2.4 (Theorem 3.2.4).
  1. $(i)$.
    The collection of tilting representations generates the category $\mathrm{D}(\mathrm{pt}/{\mathbf{G}}^{\mathrm{F}},\Lambda)$.
  2. $(ii)$.
    Tilting representations are concentrated in degree 0 and are projective representations.
These objects provide a particularly nice collection of projective generators of the category $\mathrm{D}(\mathrm{pt}/{\mathbf{G}}^{\mathrm{F}},\Lambda)$. It should be noted that we could have defined their characters previously in a formal way but the existence of $\mathcal{O}^{\mathrm{DL}}$ allows us to work with actual representations instead of their characters. The construction of these tilting representations is a key example of the benefits of working with $\mathcal{O}^{\mathrm{DL}}$ instead of uniform characters as they arise in a non-trivial way from the gluing of the various strata of $\mathcal{O}^{\mathrm{DL}}$.
1.2.3. A conjecture of Dudas–Malle
Classical category $\mathcal{O}$ is equipped with a ‘mixed refinement’ $\mathcal{O}^{\mathrm{mix}}$ and there is an involutive automorphism of $\mathcal{O}^{\mathrm{mix}}$
$\kappa:\mathcal{O}^{\mathrm{mix}}\to\mathcal{O}^{\mathrm{mix}}$
called Koszul duality which exchanges simple and tilting objects, we refer to [BGS96], [BBM04], [BY13].
In the context of representations of finite groups of Lie type there is an involutive automorphism of the representation ring $R({\mathbf{G}}^{\mathrm{F}})$ of ${\mathbf{G}}^{\mathrm{F}}$ given by Alvis–Curtis duality, see [Alv79], [Cur80], [DL82] :
$d:R({\mathbf{G}}^{\mathrm{F}})\to R({\mathbf{G}}^{\mathrm{F}}).$
It is known that in $R({\mathbf{G}}^{\mathrm{F}})$ there is an equality of characters [DL83].
$d(R_{w}^{\theta})=(-1)^{\ell(w)}R_{w}^{\theta}.$
More generally, Lusztig, in [Lus84], has given a formula for the Alvis–Curtis dual of $\mathrm{IH}(\overline{X(w)},\overline{\mathbb{Q}}_{\ell})]$. We reinterpret his calculation as follows
Theorem 1.2.5 (Corollary 4.1.4).
There is an equality
$d([\mathrm{IH}(\overline{X(w)},\overline{\mathbb{Q}}_{\ell})])=\pm d(\mathrm{ch}(T_{w}))$
where $T_{w}$ is the indecomposable tilting sheaf containing $\Delta_{w,1}$ and supported on the closure of $\mathbf{B}w\mathbf{B}$.
The calculations of Lusztig and our theorem admit refinements which take into account the weights of the Frobenius endomorphism acting on the cohomology (and intersection homology) of the Deligne–Lusztig varieties. This calculation suggests some deep interplay between Koszul duality for category $\mathcal{O}$ and Alvis–Curtis duality.
Finally, let $\delta>0$ be the smallest integer such that $(\mathbf{G},\mathrm{F}^{\delta})$ is split. Then $\mathrm{F}^{\delta}$ acts on the cohomology of $\mathrm{IH}(\overline{X(w)},\overline{\mathbb{Q}}_{\ell})]$. Let $\lambda\in\overline{\mathbb{Q}}_{\ell}$ and denote by $\mathrm{IH}(\overline{X(w)},\overline{\mathbb{Q}}_{\ell})[\lambda]]$ the character of the generalized eigenspace of $\mathrm{F}^{\delta}$ on $\mathrm{IH}(\overline{X(w)},\overline{\mathbb{Q}}_{\ell})$ for the eigenvalue $\lambda$. Let $\overline{\lambda}\in\overline{\mathbb{F}}_{\ell}$ and let $\mathrm{IH}(\overline{X(w)},\overline{\mathbb{Q}}_{\ell})[\overline{\lambda}]]$ be the character of
$\mathrm{IH}(\overline{X(w)},\overline{\mathbb{Q}}_{\ell})[\overline{\lambda}]]=\bigoplus_{\lambda}[\mathrm{IH}(\overline{X(w)},\overline{\mathbb{Q}}_{\ell})[\lambda]]$
where $\lambda$ ranges through the set of element of $\overline{\mathbb{Z}}_{\ell}$ that reduce to $\overline{\lambda}$ in $\overline{\mathbb{F}}_{\ell}$.
Theorem 1.2.6 (Theorem 4.3.1, [DM14, Conjecture 1.2]).
There exists $\ell_{0}$ depending only on $\mathbf{G}$ (not on $\mathrm{F}$) such that for all $\ell\geq\ell_{0}$, $\overline{\lambda}\in\overline{\mathbb{F}}_{\ell}$, $w\in\mathbf{W}$ the character of ${\mathbf{G}}^{\mathrm{F}}$
$d([\mathrm{IH}(\overline{X(w)},\overline{\mathbb{Q}}_{\ell})[\overline{\lambda}]])$
is the unipotent part of a projective $\overline{\mathbb{Z}}_{\ell}$-character.
Remark 1.2.7.
The hypothesis on $\ell$ in the theorem is to ensure that Kazhdan–Lusztig polynomials and $\ell$-Kazhdan–Lusztig polynomials agree. This ensure that if $T^{\overline{\mathbb{Z}}_{\ell}}_{w}$ is a $\overline{\mathbb{Z}}_{\ell}$ tilting representation, then the unipotent part of $T^{\overline{\mathbb{Z}}_{\ell}}_{w}[\frac{1}{\ell}]$ is $T_{w}^{\overline{\mathbb{Q}}_{\ell}}$ whose character is known. In general, the unipotent part of $T^{\overline{\mathbb{Z}}_{\ell}}_{w}[\frac{1}{\ell}]$ decomposes as
$T^{\overline{\mathbb{Q}}_{\ell}}_{w}\oplus_{v<w}T_{v}^{\overline{\mathbb{Q}}_{\ell},\oplus n_{w,v}}$
where $n_{w,v}\in\mathbb{N}$ are certain multiplicities.
1.2.4. Applications to decomposition numbers
The conjecture of [DM14], now theorem 4.3.1 was assumed to prove certain results concerning the unipotent decomposition numbers of ${\mathbf{G}}^{\mathrm{F}}$. Now that this conjecture is established let us mention a few consequences. Let us fix $q$ and $\ell$ such that theorem 4.3.1 holds.
  1. $(i)$.
    Assume that condition $(ii)$ of [DM14, Proposition 2.1] holds. Then the unipotent decomposition matrix of ${\mathbf{G}}^{\mathrm{F}}$ has unitriangular shape and its entries do not depend on $q$ or $\ell$. Note that the unitriangularity has also been established by other methods in [BDT20].
  2. $(ii)$.
    The conjecture of [DM14] is assumed in calculations of [DM19] and [DM15] and is used to show that certain entries vanish. This now holds under our standing assumptions on $q$ and $\ell$.
1.3. Notation and conventions
We fix $\ell\neq p$ a prime and denote by $\Lambda\in\{\overline{\mathbb{Q}}_{\ell},\overline{\mathbb{Z}}_{\ell},\overline{\mathbb{F}}_{\ell}\}$ a coefficient ring. For $X$ an algebraic stack, we denote by $\mathrm{D}(X,\Lambda)$ the category of ind-constructible sheaves of $\Lambda$-modules as defined in [HRS23]. If $X$ is a finite type scheme then we have,
$\mathrm{D}(X,\Lambda)=\mathrm{Ind}(\mathrm{D}^{b}_{c}(X,\Lambda))$
the ind-completion of the usual derived category of constructible sheaves on $X$. If the coefficient ring is clear, we will write $\mathrm{D}(X,\Lambda)=\mathrm{D}(X)$. Let $f:X\to Y$ be a morphism of stacks, whenever they are defined, we will denote by $f_{*},f_{!},f^{*},f^{!}$ the usual pushforward and pullbacks. We fix a square root of $p$ in $\overline{\mathbb{Z}}_{\ell}$. Whenever $X$ is a stack defined over $\mathbb{F}_{q}$, this defines a half Tate twist $(\frac{1}{2})$ on sheaves on $X_{\overline{\mathbb{F}}_{q}}$ with a Weil-structure. Given a stack $X$ with a Frobenius $\mathrm{F}:X\to X$, we denote by $\mathrm{D}^{b}_{c}(X)^{\mathrm{Weil}}$ the category of Weil sheaves on $X$, i.e. the category of pairs $(A,\phi)$ where $A\in\mathrm{D}^{b}_{c}(X)$ and $\phi:\mathrm{F}^{*}A\xrightarrow{\sim}A$.
We denote by $\mathrm{Hom}(-,-)$ the derived functor of endomorphism and we denote by $\mathrm{Hom}^{i}$ its $i$-th cohomology groups and by $\mathrm{End}(X)$ the functor $\mathrm{Hom}(X,X)$ (where everything is derived).
All categories are considered as $\infty$-categories but the $\infty$-part plays no essential role, all $\infty$-categories in this paper are either stable $\infty$-categories, as in [Lur17] or abelian/additive usual $1$-categories, the unfamiliar reader should ignore these technicalities and only consider triangulated categories. If $C$ is equipped with a $t$-structure, we denote by $C^{\heartsuit}$ its heart. We recall that an object $c\in C$ is compact if and only if
$\mathrm{Hom}_{C}(c,-)$
commutes with arbitrary direct sums. In particular for a ring $A$, the compact objects of $\mathrm{D}(A)$, the full derived category of $A$, are exactly the perfect complexes. They form a category which we denote by $\mathrm{Perf}(A)\subset\mathrm{D}(A)$. A complex $K\in\mathrm{D}(A)$ lies in $\mathrm{Perf}(A)$ if and only if it is quasi-isomorphic to a bounded complex of projectives of finite type.
1.4. Acknowledgments
The author thanks Olivier Dudas for suggesting the problem and for comments on a draft of this paper. We thank Jean-François Dat and Olivier Dudas for their support during the preparation of this paper. We thank Gunter Malle for doing a careful reading of a draft of this paper. This paper was written while the author was a guest at the Max Planck Institute for mathematics in Bonn. The author thanks Cédric Bonnafé and Gerhard Hiss for helpful discussions. The author thanks Xinwen Zhu for comments on a previous version of this paper.
2. Deligne–Lusztig category $\mathcal{O}$
As in the introduction, we will consider the algebraic stack
$\frac{\mathbf{U}\backslash\mathbf{G}/\mathbf{U}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{T}}$
which we call the horocycle stack. We denote by $\mathcal{O}^{\mathrm{DL}}_{\Lambda}$ the category $\mathrm{D}(\frac{\mathbf{U}\backslash\mathbf{G}/\mathbf{U}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{T}},\Lambda)$. We call this category the Deligne–Lusztig category $\mathcal{O}$. The goal of this section is to discuss the homological structure of this category. When the coefficients are clear, we will drop the index $\Lambda$ from the notation. We denote by $\mathcal{O}_{\mathrm{c}}^{\mathrm{DL}}$ the full subcategory of constructible sheaves.
2.1. Stratification
We equip the horocycle stack with its Bruhat stratification, that is
$\frac{\mathbf{U}\backslash\mathbf{G}/\mathbf{U}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{T}}=\bigsqcup_{w\in\mathbf{W}}\frac{\mathbf{U}\backslash\mathbf{B}w\mathbf{B}/\mathbf{U}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{T}}.$
We denote by $j_{w}:\frac{\mathbf{U}\backslash\mathbf{B}w\mathbf{B}/\mathbf{U}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{T}}\to\frac{\mathbf{U}\backslash\mathbf{G}/\mathbf{U}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{T}}$ the inclusion of the stratum $w$, note that this is an affine immersion.
Lemma 2.1.1.
Let $\dot{w}\in\mathbf{N}(\mathbf{T})$ be a lift of $w$. There is an isomorphism of stacks
$\frac{\mathbf{U}\backslash\mathbf{B}w\mathbf{B}/\mathbf{U}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{T}}=\mathrm{pt}/({\mathbf{T}}^{w\mathrm{F}}\ltimes\mathbf{U}_{w})$
where $\mathbf{U}_{w}=\mathbf{U}\cap\mathrm{Ad}(\dot{w})(\mathbf{U})$.
Proof.
See [Ete24, Lemma 3.2.4]. ∎
Since $\mathbf{U}_{w}$ is a connected unipotent group, there is an equivalence of categories
$\mathrm{D}(\mathrm{pt}/({\mathbf{T}}^{w\mathrm{F}}\rtimes\mathbf{U}_{w}))=\mathrm{D}(\mathrm{Rep}_{\Lambda}{\mathbf{T}}^{w\mathrm{F}})$
see [Ete24, Section 3.2] for a discussion on the normalization of this equivalence. This category decomposes into blocks
$\mathrm{D}(\mathrm{Rep}_{\Lambda}{\mathbf{T}}^{w\mathrm{F}})=\oplus_{\chi}\mathrm{D}^{\chi}(\mathrm{Rep}_{\Lambda}{\mathbf{T}}^{w\mathrm{F}})$
where $\chi$ ranges through the irreducible characters of ${\mathbf{T}}^{w\mathrm{F}}$ over $\overline{\mathbb{Q}}_{\ell}$ if $\Lambda=\overline{\mathbb{Q}}_{\ell}$ and over $\overline{\mathbb{F}}_{\ell}$ is $\Lambda=\overline{\mathbb{F}}_{\ell},\overline{\mathbb{Z}}_{\ell}$. The unipotent block (with respect to $\Lambda$) is the block containing the trivial representation. This is usually also called the principal block.
Let $\rho$ be a representation of ${\mathbf{T}}^{w\mathrm{F}}$, we denote by
  1. $(i)$.
    $\Delta_{w}(\rho)=j_{w,!}\rho[\ell(w)](\frac{\ell(w)}{2})$,
  2. $(ii)$.
    $\nabla_{w}(\rho)=j_{w,*}\rho[\ell(w)](\frac{\ell(w)}{2})$.
Since the $j_{w}$ are affine immersions, these sheaves are perverse sheaves in $\mathcal{O}^{\mathrm{DL}}$. We have indicated the Tate twist in this definition in anticipation of section 2.5, they will not play a role before this section.
For a character $\chi:{\mathbf{T}}^{w\mathrm{F}}\to\Lambda^{\times}$ of ${\mathbf{T}}^{w\mathrm{F}}$, we denote by $E_{\chi}$ the projective cover of $\chi$ and by
  1. $(i)$.
    $\Delta_{w,\chi}=\Delta_{w}(E_{\chi})$,
  2. $(ii)$.
    $\nabla_{w,\chi}=\nabla_{w}(E_{\chi})$.
The objects $\Delta_{w,\chi}$ and $\nabla_{w,\chi}$ are called the standard and costandard objects respectively. Note that if $\Lambda=\overline{\mathbb{Q}}_{\ell}$, then $\chi=E_{\chi}$.
Finally, we will also denote for a representation $\rho$ of ${\mathbf{T}}^{w\mathrm{F}}$
$\mathrm{IC}_{w}(\rho)=j_{w,!*}\rho[\ell(w)](\frac{\ell(w)}{2}).$
Proposition 2.1.2.
  1. $(i)$.
    The category $\mathcal{O}^{\mathrm{DL},\heartsuit}$ is noetherian.
  2. $(ii)$.
    If $\Lambda$ is a field, it is also Artinian and has finitely many irreducible objects which are the $\mathrm{IC}_{w}(\chi)$ where $\chi$ ranges through the set of irreducible representations of ${\mathbf{T}}^{w\mathrm{F}}$.
Proof.
This is a direct application of [BBDG18, Theorem 4.3.1]. ∎
Proposition 2.1.3.
The category $\mathcal{O}^{\mathrm{DL}}$ is compactly generated. The compact objects are exactly the sheaves that are constructible and compact when *-pullbacked (equivalently !!-pullbacked) to all strata. In particular, the standard and costandard objects are compact.
We denote by $\mathcal{O}^{\mathrm{DL},\omega}$ the full subcategory of compact objects.
Remark 2.1.4.
The reader should note that if $\Lambda\in\{\overline{\mathbb{Z}}_{\ell},\overline{\mathbb{F}}_{\ell}\}$ then the inclusion $\mathcal{O}^{\mathrm{DL},\omega}\subset\mathcal{O}^{\mathrm{DL}}_{\mathrm{c}}$ can be strict. This already happens in the case when $G$ is a torus as the horocycle stack is then $\mathrm{pt}/{\mathbf{T}}^{\mathrm{F}}$ and the category of sheaves on it is simply the category of representations of ${\mathbf{T}}^{\mathrm{F}}$. The constructible sheaves on ${\mathbf{T}}^{\mathrm{F}}$ are exactly the complexes of representations whose underlying complex of $\Lambda$-modules is perfect but if $\ell||{\mathbf{T}}^{\mathrm{F}}|$ then the trivial representation has infinite cohomological dimension and is therefore constructible but not compact.
Remark 2.1.5.
It follows from Proposition 2.1.3 that if $\Lambda=\overline{\mathbb{Q}}_{\ell}$ then the inclusion $\mathcal{O}^{\mathrm{DL},\omega}\subset\mathcal{O}_{\mathrm{c}}^{\mathrm{DL}}$ is an equivalence.
Proof of proposition 2.1.3.
The category $\mathcal{O}^{\mathrm{DL}}$ is generated by the objects $\Delta_{w,\chi}$. Since the immersions $j_{w}$ are of finite presentation both functors $j_{w}^{!}$ and $j_{w,*}$ are continuous hence their left adjoints preserve compact objects. If follows that the standard sheaves form a set of compact generators of the category.
Let $A$ be a compact object. By the previous point $j_{w}^{*}A$ is compact and since in $\mathrm{D}(\mathrm{Rep}_{\Lambda}{\mathbf{T}}^{w\mathrm{F}})$ the compact objects are stable under Verdier duality, we deduce that $j_{w}^{!}A$ is compact. Conversely let $A$ be such that for all $w$, the objects $j_{w}^{!}A$ and $j_{w}^{*}A$ are compact. Let $w$ be such that $A$ is supported on $\mathbf{B}w\mathbf{B}$ and this stratum is maximal for the inclusion and consider $C=\mathrm{cone}(j_{w,!}j_{w}^{!}A\to A)$ the cone of the adjunction map. Let $i_{w}$ be the inclusion of the union of the strata in the closure of $\mathbf{B}w\mathbf{B}$, then $i_{w}$ is a closed immersion of finite presentation and therefore $i_{w}^{*}$ and $i_{w,*}$ both preserve compact objects, hence $C=i_{w,*}i_{w}^{*}A$ is compact.
It remains to show that the costandard objects are compact. It is enough to show that $j_{v}^{*}\nabla_{w,\chi}$ is compact which follows from applying Verdier duality and the case of the standard sheaves. ∎
2.2. Link with Hecke categories
Hecke categories attached to $\mathbf{G}$ have two realizations : the equivariant Hecke categories and the free monodromic Hecke categories. These categories were studied, in particular, in [BGS96], [LY20], [BY13], [BR22], [Gou22], [Ete23]. Let us give a brief account of these categories. We denote by $\pi_{1}^{t}(\mathbf{T})$ the tame quotient of the étale fundamental group of the torus $\mathbf{T}$ at the geometric point $1\in\mathbf{T}$. It is known that there is an isomorphism
$\pi_{1}^{t}(\mathbf{T})=X_{*}(\mathbf{T})\otimes\pi_{1}^{t}(\mathbb{G}_{m})=X_{*}(\mathbf{T})\otimes\hat{\mathbb{Z}}^{(p)}(1).$
Let $\chi,\chi^{\prime}:\pi_{1}^{t}(\mathbf{T})\to\Lambda^{\times}$ be two characters of finite order and let us denote by $\mathcal{L}_{\chi}$ (reps. $\mathcal{L}_{\chi^{\prime}}$) the corresponding Kummer sheaves on $\mathbf{T}$. We can also pull them back to $\mathbf{B}$ and understand them as sheaves on $\mathbf{B}$. With this data, one can construct two categories :
  1. $(i)$.
    the category $\mathrm{D}^{b}_{c}((\mathbf{B},\mathcal{L}_{\chi})\backslash\mathbf{G}/(\mathbf{B},\mathcal{L}_{\chi^{\prime}}))$ of $(\mathbf{B}\times\mathbf{B},\mathcal{L}_{\chi}\otimes\mathcal{L}_{\chi^{\prime}})$-equivariant sheaves on $\mathbf{G}$, we refer to [LY20] for their construction and properties,
  2. $(ii)$.
[…]
In all the previous papers, the equivariant (resp. monodromic) Hecke category is the category
[…]
We now introduce functors between these Hecke categories and $\mathcal{O}^{\mathrm{DL}}$.
  1. $(i)$.
    The free monodromic Hecke category is realized as a category of sheaves on $\mathbf{U}\backslash\mathbf{G}/\mathbf{U}$, let $\mathfrak{p}:\mathbf{U}\backslash\mathbf{G}/\mathbf{U}\to\frac{\mathbf{U}\backslash\mathbf{G}/\mathbf{U}}{\mathrm{Ad}_{\mathrm{F}}(\mathbf{T})}$ be the natural quotient map. We will consider $\mathfrak{p}_{!}:\mathcal{H}^{\mathrm{mon}}\to\mathcal{O}^{\mathrm{DL}}$,
  2. $(ii)$.
    Let $(\chi,\chi^{\prime})$ be such that there exists an element $w\in\mathbf{W}$ such that $\chi^{\prime}=w\mathrm{F}(\chi)$, then there is a natural forgetful functor $\mathfrak{q}^{*}:\mathrm{D}^{b}_{c}((\mathbf{B},\mathcal{L}_{\chi})\backslash\mathbf{G}/(\mathbf{B},\mathcal{L}_{\chi^{\prime}}))\to\mathcal{O}^{\mathrm{DL}}$ which is obtained by forgetting the $(\mathbf{T}\times\mathbf{T},\mathcal{L}_{\chi}\otimes\mathcal{L}_{\chi^{\prime}})$-equivariance down to $\mathbf{T}$-equivariance along the inclusion $\mathbf{T}\to\mathbf{T}\times\mathbf{T}$ given by $t\mapsto(t,\mathrm{F}(t^{-1}))$. We denote it by $\mathfrak{q}^{*}$ as we think of this functor as a pullback along some map $\frac{\mathbf{U}\backslash\mathbf{G}/\mathbf{U}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{T}}\to(\mathbf{B},\mathcal{L}_{\chi})\backslash\mathbf{G}/(\mathbf{B},\mathcal{L}_{\chi^{\prime}})$.
Lemma 2.2.1.
Let $w\in\mathbf{W}$ and $\chi$ be a character of $\mathbf{T}^{w\mathrm{F}}$. There are natural isomorphisms
  1. $(i)$.
    $\mathfrak{p}_{!}\Delta_{w,\chi}^{\mathrm{mon}}[\dim\mathbf{T}](\frac{\dim\mathbf{T}}{2})=\Delta_{w}(E_{\chi}),\mathfrak{p}_{!}\nabla_{w,\chi}^{\mathrm{mon}}[\dim\mathbf{T}](\frac{\dim\mathbf{T}}{2})=\nabla_{w}(E_{\chi})$
  2. $(ii)$.
    $\mathfrak{q}^{*}\Delta_{w,\chi}^{\mathrm{eq}}=\Delta_{w}(\chi),\mathfrak{q}^{*}\nabla_{w,\chi}^{\mathrm{eq}}=\nabla_{w}(\chi)$.
Proof.
The second point is clear since $\mathfrak{q}^{*}$ is simply the forgetful functor. The first point follows from [Ete24, Lemma 2.8.3]. ∎
2.3. Tilting objects
Definition 2.3.1.
Let $A\in\mathcal{O}^{\mathrm{DL},\heartsuit}$,
  1. $(i)$.
    a $\Delta$-filtration on $A$ is a filtration whose graded pieces are standard sheaves,
  2. $(ii)$.
    a $\nabla$-filtration on $A$ is a filtration whose graded pieces are costandard sheaves,
  3. $(iii)$.
    the sheaf $A$ is tilting if $A$ has both a $\Delta$ and a $\nabla$-filtration.
We denote by $\mathcal{O}^{\mathrm{DL}}_{\mathrm{tilt}}$ the full subcategory of $\mathcal{O}^{\mathrm{DL},\heartsuit}$ of tilting objects.
Theorem 2.3.2.
  1. $(i)$.
    The category $\mathcal{O}^{\mathrm{DL},\heartsuit}$ has enough projective objects and all projectives have a $\Delta$-filtration. If $\Lambda$ is a field it also has enough injective objects and the injective objects have a $\nabla$-filtration.
  2. $(ii)$.
    The category $\mathcal{O}^{\mathrm{DL}}_{\mathrm{tilt}}$ is an additive, idempotent complete and Krull-Schmidt category.
  3. $(iii)$.
    There is a unique bijection between indecomposable tilting objects and pairs $(w,\chi)$ where $w\in\mathbf{W}$ and $\chi$ is a irreducible character (over the residue field of $\Lambda$) of $\mathbf{T}^{w\mathrm{F}}$ characterized by the fact that the indecomposable tilting corresponding to $(w,\chi)$ is supported on the closure of $\mathbf{B}w\mathbf{B}$ and has $\Delta_{w,\chi}$ with multiplicity one in a $\Delta$-filtration. We denote this tilting object by $T_{w,\chi}$.
Remark 2.3.3.
Theorem 2.3.2 is very analogous to the structure theorem for category $\mathcal{O}$, see [BGS96, Section 3] and the proof uses the yoga of highest weight categories, see [Ric16, Section 7] and [AR16] for the $\mathbb{Z}_{\ell}$-case.
Proof of Theorem 2.3.2.
The existence of enough projective for $\mathcal{O}^{\mathrm{DL},\heartsuit}$ follows from the same statement established for free monodromic Hecke categories and the next two observations.
  1. $(i)$.
    If $P$ is a free monodromic projective object then $\mathfrak{p}_{!}P[\dim\mathbf{T}]$ is perverse. Indeed free monodromic projective objects have a $\Delta^{\mathrm{mon}}$-filtration, see [BY13] and [Gou22], hence $\mathfrak{p}_{!}P[\dim\mathbf{T}]$ has a $\Delta$-filtration by lemma 2.2.1.
  2. $(ii)$.
    The right adjoint functor $\mathfrak{p}^{!}[-\dim\mathbf{T}]$ is $t$-exact, as $\mathfrak{p}$ is smooth, hence $\mathrm{Hom}(\mathfrak{p}_{!}P,-)[-\dim\mathbf{T}]=\mathrm{Hom}(P,\mathfrak{p}^{!}-)[-\dim\mathbf{T}]$ is an exact functor.
The existence of a $\Delta$-filtration is therefore established for a generating family of projective objects. The existence for all projective objects follows from the characterization of objects with a $\Delta$-filtration of Lemma 2.3.5 and the fact that this characterization is stable under taking direct summands. The statement about injective objects follows by Verdier duality. This establishes $(i)$.
For $(ii)$, Lemma 2.3.5 guarantees that the category $\mathcal{O}^{\mathrm{DL}}_{\mathrm{tilt}}$ is idempotent complete and additive. The fact that it is Krull-Schmidt follows from the constructibility of tilting objects and the fact that $\Lambda$ is a local ring.
For the characterization of indecomposable tilting objects, we first note that the functor $\mathfrak{p}_{!}[\dim\mathbf{T}]$ sends tilting objects to tilting objects by Lemma 2.2.1. For free monodromic Hecke categories, the corresponding statement is established in [BY13] in the unipotent case and [Gou22] in the non-unipotent case. This guarantees the existence of a tilting sheaf supported on the closure of $\mathbf{B}w\mathbf{B}$ and such that the multiplicity of $\Delta_{w,\chi}$ in one of its $\Delta$-flags is $1$. Since the category of tilting sheaves is Krull-Schmidt there exists a direct summand that is indecomposable and satisfies both of these properties. The uniqueness can be proved as in [Ric16, Theorem 7.14]. ∎
Remark 2.3.4.
Note that we prove the existence of a generating collection of tilting sheaves from the existence of tilting sheaves in the Hecke category. Also note that the pushforward of an indecomposable tilting sheaf is indecomposable. Indeed if $T\in\mathcal{H}^{\mathrm{mon}}$ is indecomposable then $\mathrm{End}(T)$ is a local algebra and $\mathrm{End}(\mathfrak{p}_{!}T)$ is a quotient of it hence it is also local.
Lemma 2.3.5.
A sheaf $A\in\mathcal{O}^{\mathrm{DL}}$ is tilting if and only if for all $w\in\mathbf{W}$, the objects $j_{w}^{!}A$ and $j_{w}^{*}A$ are perverse and projective representations of $\mathbf{T}^{w\mathrm{F}}$. More generally $A$ has a $\Delta$ (reps. $\nabla$)-filtration if and only if $j_{w}^{*}A$ is perverse and projective (resp. $j_{w}^{!}A$ is perverse and projective).
Proof.
If $A$ has a $\Delta$-filtration, then $j_{w}^{*}A$ is filtered with graded pieces of the form $j_{w}^{*}\Delta_{v,\chi}$, where $v\in\mathbf{W}$ and $\chi$ is a character of $\mathbf{T}^{v\mathrm{F}}$. Since we have
  1. $(i)$.
    $j_{w}^{*}\Delta_{v,\chi}=0$ if $v\neq w$,
  2. $(ii)$.
    $j_{w}^{*}\Delta_{v,\chi}=E_{\chi}$ if $v=w$,
it follows that $j_{w}^{*}$ is perverse and projective. Conversely if for all $w$, the sheaf $j_{w}^{*}A$ is perverse and projective, we now show that it is equipped with a $\Delta$-filtration. We equip $\mathbf{W}$ with the Bruhat order, i.e. the order induced by the closure relations of strata and we choose an extension of this order to a total order, we denote this extension $\leq_{\mathrm{tot}}$. Let $w$ be the minimal element of $\mathbf{W}$ such that $j_{w}^{*}A\neq 0$. We denote by $i_{w}:\frac{\mathbf{U}\backslash\overline{\mathbf{B}w\mathbf{B}}/\mathbf{U}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{T}}\subset\frac{\mathbf{U}\backslash\mathbf{G}/\mathbf{U}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{T}}$ the inclusion of the closure of the stratum corresponding to $w$ and by $k_{w}:V_{w}\subset\frac{\mathbf{U}\backslash\mathbf{G}/\mathbf{U}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{T}}$ the open complement. Since $w$ is minimal among the strata supporting $A$, we have
$j_{w,!}j_{w}^{*}A=i_{w,!}i_{w}^{*}A,$
which is a direct sum of standard sheaves. We now have an excision triangle
$k_{w,!}k_{w}^{*}A\to A\to i_{w,!}i_{w}^{*}A.$
The map $A\to i_{w,!}i_{w}^{*}A$ is a map of perverse sheaves that is surjective in the category of perverse sheaves. Indeed the cokernel of this map is supported on $\frac{\mathbf{U}\backslash{\mathbf{B}w\mathbf{B}}/\mathbf{U}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{T}}$ but the map $A\to i_{w,!}i_{w}^{*}A$ is an isomorphism when restricted to $\frac{\mathbf{U}\backslash{\mathbf{B}w\mathbf{B}}/\mathbf{U}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{T}}$. It follows that this triangle is a short exact sequence of perverse sheaves. Finally, the sheaf $B=k_{w,!}k_{w}^{*}A$ satisfies
  1. $(i)$.
    $j_{v}^{*}B=j_{v}^{*}A$ if $v>_{\mathrm{tot}}w$ and is perverse and projective by induction,
  2. $(ii)$.
    $j_{v}^{*}B=0$ if $v\leq_{\mathrm{tot}}w$.
Since $A$ is supported on finitely many strata (as $\mathbf{W}$ is finite), by induction we can reduce to the case where $A$ is supported on a single stratum but then the claim is trivial. ∎
Lemma 2.3.6.
Let $T,T^{\prime}$ be two tilting sheaves then $\mathrm{Hom}(T,T^{\prime})$ is concentrated in degree 0.
Proof.
By the five lemma, this reduces to the fact that $\mathrm{Hom}(\Delta,\nabla)$ is concentrated in degree 0. ∎
Proposition 2.3.7.
The category $\mathcal{O}^{\mathrm{DL},\omega}$ is generated by $\mathcal{O}^{\mathrm{DL}}_{\mathrm{tilt}}$ as a triangulated category.
Proof.
This is clear since the standard and tilting objects generate the same category. ∎
Lemma 2.3.8.
Let $T\in\mathcal{O}_{\mathrm{tilt},\overline{\mathbb{Z}}_{\ell}}^{\mathrm{DL}}$ then $T\otimes_{\overline{\mathbb{Z}}_{\ell}}\overline{\mathbb{Q}}_{\ell}$ and $T\otimes_{\overline{\mathbb{Z}}_{\ell}}\overline{\mathbb{F}}_{\ell}$ are tilting. Moreover $T$ is indecomposable if and only if $T\otimes_{\overline{\mathbb{Z}}_{\ell}}\overline{\mathbb{F}}_{\ell}$ is indecomposable.
Proof.
The object $T$ is indecomposable if and only if $\mathrm{End}(T)$ is a local algebra. As $T$ is tilting $\mathrm{End}(T)$ is a finite free $\overline{\mathbb{Z}}_{\ell}$-module and the natural map $\mathrm{End}(T)\otimes_{\overline{\mathbb{Z}}_{\ell}}\overline{\mathbb{F}}_{\ell}\to\mathrm{End}(T\otimes_{\overline{\mathbb{Z}}_{\ell}}\overline{\mathbb{F}}_{\ell})$ is an isomorphism. Hence $\mathrm{End}(T)$ is local if and only $\mathrm{End}(T)\otimes_{\overline{\mathbb{Z}}_{\ell}}\overline{\mathbb{F}}_{\ell}$ is local. ∎
2.4. Action of $\mathbf{T}^{\vee}$
Consider the action of $\mathbf{T}\times\mathbf{T}$ on the horocycle stack by left and right translations. We denote by $\mathbf{T}^{\vee}$ the torus dual to $\mathbf{T}$ defined over $\Lambda$ and by $\mathrm{F}^{\vee}:\mathbf{T}^{\vee}\to\mathbf{T}^{\vee}$ the morphism dual to $\mathrm{F}$.
Lemma 2.4.1.
All sheaves on $\frac{\mathbf{U}\backslash\mathbf{G}/\mathbf{U}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{T}}$ are $\mathbf{T}\times\mathbf{T}$-monodromic. Moreover the monodromy map yields a $(\mathbf{T}^{\vee}\sslash\mathbf{W})^{\mathrm{F}^{\vee}}$-linear structure on $\mathcal{O}^{\mathrm{DL}}$.
Proof.
Recall that monodromic sheaves are sheaves that are locally constant along $\mathbf{T}\times\mathbf{T}$-orbits and tame local systems, we refer to [Ete24, Section 2] for a discussion. Since all strata are classifying spaces the monodromic part is clear. Monodromic sheaves are equipped with a canonical action of $\pi_{1}^{t}(\mathbf{T}\times\mathbf{T})$, the tame fundamental group of $\mathbf{T}\times\mathbf{T}$. This groups is isomorphic to $X_{*}(\mathbf{T}\times\mathbf{T})\otimes\hat{\mathbb{Z}}^{(p)}(1)$. After fixing a trivialization of the roots of $1$ in $\overline{\mathbb{F}}_{q}$ we get a topological generator of $\hat{\mathbb{Z}}^{(p)}(1)$ and an action of $\Lambda[X_{*}(\mathbf{T}\times\mathbf{T})]=\mathcal{O}(\mathbf{T}^{\vee})\otimes_{\Lambda}\mathcal{O}(\mathbf{T}^{\vee})$ on monodromic sheaves. We first restrict this action to an action of $\mathcal{O}(\mathbf{T}^{\vee})^{\mathbf{W}}\otimes_{\Lambda}\mathcal{O}(\mathbf{T}^{\vee})^{\mathbf{W}}$. Finally we observe that the $\mathrm{Ad}_{\mathrm{F}}(\mathbf{T})$-equivariance shows that the map
$\mathcal{O}(\mathbf{T}^{\vee})^{\mathbf{W}}\otimes_{\Lambda}\mathcal{O}(\mathbf{T}^{\vee})^{\mathbf{W}}\to\mathrm{maps}(\mathrm{id}_{\mathcal{O}^{\mathrm{DL}}})$
factors through the quotient of $\mathcal{O}({\mathbf{T}}^{\vee})^{\mathbf{W}}\otimes_{\Lambda}\mathcal{O}({\mathbf{T}}^{\vee})^{\mathbf{W}}$ by the ideal $(f-\mathrm{F}^{\vee}(f),f\in\mathcal{O}((\mathbf{T}^{\vee})^{\mathbf{W}})$, that is, by $\mathcal{O}(\mathbf{T}^{\vee}\sslash\mathbf{W})^{\mathrm{F}^{\vee}})$. ∎
Remark 2.4.2.
It should be noted that there is a bigger scheme acting on $\mathcal{O}^{\mathrm{DL}}$. However, this finer structure will not be compatible when we go to $\mathbf{G}^{\mathrm{F}}$-representations.
Lemma 2.4.3.
There is a bijection, depending on a choice of a trivialization of the roots of $1$ in $\overline{\mathbb{F}}_{q}$, between $\Lambda$-points of $(\mathbf{T}^{\vee}\sslash\mathbf{W})^{\mathrm{F}^{\vee}}$ and $\mathrm{F}^{*}$-stable semisimple $\mathbf{G}^{*}$-conjugacy classes of elements of order invertible in $\Lambda$. We denote this set $\mathbf{G}^{*}_{\Lambda}/\mathrm{ss}$.
Proof.
Both sets are identified with the set of $\mathbf{W}$-conjugacy classes of characters $X_{*}(\mathbf{T})\to\Lambda^{\times}$ that are stable under the Frobenius, see [DL76, Corollary 5.24]. ∎
Corollary 2.4.4.
The category $\mathcal{O}^{\mathrm{DL}}$ splits as
$\mathcal{O}^{\mathrm{DL}}=\bigoplus_{s\in(\mathbf{G}^{*}_{\Lambda}/\mathrm{ss})^{\mathrm{F}^{*}}}\mathcal{O}^{\mathrm{DL},s}$
where $\mathcal{O}^{\mathrm{DL},s}$ is the full subcategory of $\mathcal{O}^{\mathrm{DL}}$ generated by $\Delta_{w,\chi}$ where the geometric conjugacy class of $(w,\chi)$ corresponds to $s$ (in the sense of [DL76, Section 5]).
Proof.
This follows immediately from Lemma 2.4.3 as the corresponding points index the connected components of $(\mathbf{T}^{\vee}\sslash\mathbf{W})^{\mathrm{F}^{\vee}}$. ∎
Remark 2.4.5.
In the following, we will denote by $\mathcal{O}^{\mathrm{DL},\mathrm{unip}}$ the direct summand corresponding to $s=1$.
Lemma 2.4.6.
Let $A\in\mathcal{O}^{\mathrm{DL},\mathrm{unip}}_{\overline{\mathbb{Z}}_{\ell}}$ then $A[\frac{1}{\ell}]=\oplus_{s_{\ell}}A_{s_{\ell}}$ where $s_{\ell}$ ranges through the set of conjugacy classes of element of order $\ell$. In particular we have
  1. $(i)$.
    $\Delta_{w,1}^{\overline{\mathbb{Z}}_{\ell}}[\frac{1}{\ell}]=\bigoplus_{\chi_{\ell}}\Delta_{w,\chi_{\ell}}^{\overline{\mathbb{Q}}_{\ell}}$,
  2. $(ii)$.
    $\nabla_{w,1}^{\overline{\mathbb{Z}}_{\ell}}[\frac{1}{\ell}]=\bigoplus_{\chi_{\ell}}\nabla_{w,\chi_{\ell}}^{\overline{\mathbb{Q}}_{\ell}}$,
  3. $(iii)$.
    $T_{w,1}^{\overline{\mathbb{Z}}_{\ell}}[\frac{1}{\ell}]=\bigoplus_{\chi_{\ell}}T_{w,\chi_{\ell}}^{\overline{\mathbb{Q}}_{\ell}}\oplus\bigoplus_{v<w,\chi_{\ell}}T_{v,\chi_{\ell}}^{n_{v,w,\chi_{\ell}}}$, where $n_{v,w,\chi_{\ell}}$ is some multiplicity.
In the previous sums the character $\chi_{\ell}$ ranges over characters of $\ell^{\infty}$-torsion of $\mathbf{T}^{w\mathrm{F}}$.
Proof.
The first point follows from the comparison between the $\overline{\mathbb{Z}}_{\ell}$ and $\overline{\mathbb{Q}}_{\ell}$-versions of Lemma 2.4.3 and Corollary 2.4.4. The decompositions of $\Delta_{w,1}^{\overline{\mathbb{Z}}_{\ell}}$ and $\nabla_{w,1}^{\overline{\mathbb{Z}}_{\ell}}$ follows from the compatibility of the formation of the functors $j_{w,!}$ and $j_{w,*}$ with change of coefficients. The decomposition of the tilting objects follows from the first observation that $T_{w,1}^{\overline{\mathbb{Z}}_{\ell}}[\frac{1}{\ell}]$ is tilting by the decomposition of the $\Delta$ and $\nabla$. Furthermore $T_{w,1}^{\overline{\mathbb{Z}}_{\ell}}[\frac{1}{\ell}]$ is supported on the closure of $\mathbf{B}w\mathbf{B}$ and contains each of the $\Delta_{w,\chi_{\ell}}^{\overline{\mathbb{Q}}_{\ell}}$ with multiplicity one, the rest of the statement follows from the classification Theorem 2.3.2. ∎
2.5. Mixed Deligne–Lusztig category $\mathcal{O}$
We equip the stack $\frac{\mathbf{U}\backslash\mathbf{G}/\mathbf{U}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{T}}$ with the endomorphism $\mathrm{F}^{\delta}$ where $\delta$ is minimal such that the pair $(\mathbf{G},\mathrm{F}^{\delta})$ is a split reductive group over $\mathbb{F}_{q^{\delta}}$. All mixed structures are done with respect to this Frobenius. Furthermore, the Bruhat stratification is stable under $\mathrm{F}^{\delta}$. We define the category $\mathcal{O}^{\mathrm{DL},\mathrm{mix}}$ to be the full subcategory of constructible mixed $\overline{\mathbb{Q}}_{\ell}$-sheaves on $\frac{\mathbf{U}\backslash\mathbf{G}/\mathbf{U}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{T}}$ generated as a triangulated category by Tate twists of pure $\mathrm{IC}$-sheaves of weight 0. There is a natural forgetful functor
$\omega:\mathcal{O}^{\mathrm{DL},\mathrm{mix}}\to\mathcal{O}_{\mathrm{c}}^{\mathrm{DL}}.$
Lemma 2.5.1.
For all pairs $(w,\chi)$ where $w\in\mathbf{W}$ and $\chi$ is a character of $\mathbf{T}^{w\mathrm{F}}$, there exist objects $\Delta_{w,\chi}^{\mathrm{mix}},\nabla_{w,\chi}^{\mathrm{mix}},\mathrm{IC}_{w,\chi}^{\mathrm{mix}}$ and $T_{w,\chi}^{\mathrm{mix}}$ whose images under $\omega$ are respectively $\Delta_{w,\chi},\nabla_{w,\chi},\mathrm{IC}_{w,\chi}$ and $T_{w,\chi}$.
Proof.
The statement for the standard and costandard sheaves follows from the fact that the functors $j_{w,!}$ and $j_{w,*}$ preserve mixed sheaves. The statement for the $\mathrm{IC}$-sheaves is clear by [BBDG18]. Finally the statement for the tilting follows from [BY13], as the tilting sheaves are obtained by pushing along $\mathbf{U}\backslash\mathbf{G}/\mathbf{U}\to\frac{\mathbf{U}\backslash\mathbf{G}/\mathbf{U}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{T}}$ the free monodromic tilting sheaves which are mixed. ∎
Remark 2.5.2.
The mixed refinements in Lemma 2.5.1 are uniquely characterized up to isomorphism by the fact that all four collections of objects are perverse, the sheaf $\mathrm{IC}_{w,\chi}^{\mathrm{mix}}$ is a simple mixed perverse sheaf and the object $T_{w,\chi}^{\mathrm{mix}}$ is a mixed tilting sheaf.
Let $\mathcal{H}$ denote the generic Hecke algebra of $\mathbf{W}$ over $\mathbb{Z}[v^{\pm 1}]$. This is the algebra generated by elements $(T_{w})$ subject to the relations
  1. $(i)$.
    $T_{w}T_{w^{\prime}}=T_{ww^{\prime}}$ if $\ell(ww^{\prime})=\ell(w)+\ell(w^{\prime})$,
  2. $(ii)$.
    $T_{s}^{2}=v^{-2}T_{e}+(v^{-2}-1)T_{s}$ if $s\in\mathbf{W}$ is a simple reflection.
We denote by $H_{w}=v^{\ell(w)}T_{w}$ and we use the notation of [Soe97]. Recall from [KL79] that there is an anti-involution
$\overline{(-)}:\mathcal{H}\to\mathcal{H}$
satisfying $\overline{v}=v^{-1}$ and $\overline{H_{w}}=H^{-1}_{w^{-1}}$. There are now two bases of $\mathcal{H}$ that are self dual, they are denoted by $(C_{w})$ and $(C^{\prime}_{w})$ in loc. cit.. We shall use the notation of [Soe97].
Theorem 2.5.3 ([KL79]).
For all $w\in\mathbf{W}$, there exist two unique self dual elements $\underline{H}_{w}$ and $\underline{\tilde{H}}_{w}$ of $\mathcal{H}$ such that
  1. $(i)$.
    $\underline{H}_{w}\in H_{w}+\sum_{y<w}v\mathbb{Z}[v]H_{y}$,
  2. $(ii)$.
    $\underline{\tilde{H}}_{w}\in H_{w}+\sum_{y<w}v^{-1}\mathbb{Z}[v^{-1}]H_{y}$.
Remark 2.5.4 ([Soe97, Remark 2.4]).
We denote by $b:\mathcal{H}\to\mathcal{H}$ the involution defined by
$b(H_{w})=H_{w},b(v)=-v^{-1}.$
Then there is an equality
$b(\underline{H}_{w})=\underline{\tilde{H}}_{w}.$
Lemma 2.5.5.
There is an isomorphism of abelian groups
$\mathcal{H}\simeq K_{0}(\mathcal{O}^{\mathrm{DL},\mathrm{mix},\mathrm{unip}}_{\overline{\mathbb{Q}}_{\ell}}).$
Moreover the following holds
$\Delta_{w,1}^{\mathrm{mix}}]=H_{w},[\nabla_{w,1}^{\mathrm{mix}}]=H_{w^{-1}}^{-1},[T_{w,1}^{\mathrm{mix}}]=\underline{H}_{w},[\mathrm{IC}_{w,1}^{\mathrm{mix}}]=\underline{\tilde{H}}_{w},[-(\frac{1}{2})]=v^{-1}[-].$
Proof.
Since $\mathcal{O}^{\mathrm{DL},\mathrm{mix},\mathrm{unip}}$ is stratified by the categories $\mathrm{D}(\mathrm{pt}/\mathbf{T}^{w\mathrm{F}})^{\mathrm{unip},\mathrm{mix}}\simeq\mathrm{D}(\mathrm{pt})^{\mathrm{mix}}$, we get an isomorphism
$\mathcal{H}\simeq K_{0}(\mathcal{O}^{\mathrm{DL},\mathrm{mix},\mathrm{unip}}_{\overline{\mathbb{Q}}_{\ell}})$
satisfying $\Delta_{w,1}^{\mathrm{mix}}]=H_{w}$. Similarly, there are isomorphisms between $\mathcal{H}$ and the Grothendieck groups of the two mixed realizations of the Hecke categories $\mathcal{H}^{\mathrm{eq},\mathrm{unip},\mathrm{mix}}$ and $\mathcal{H}^{\mathrm{mon},\mathrm{unip},\mathrm{mix}}$ constructed in [BY13], that is we have
$K_{0}(\mathcal{H}^{\mathrm{eq},\mathrm{unip},\mathrm{mix}})=\mathcal{H}=K_{0}(\mathcal{H}^{\mathrm{mon},\mathrm{unip},\mathrm{mix}}).$
The quotient maps
$\mathbf{U}\backslash\mathbf{G}/\mathbf{U}\to\frac{\mathbf{U}\backslash\mathbf{G}/\mathbf{U}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{T}}\to\mathbf{B}\backslash\mathbf{G}/\mathbf{B}$
yield isomorphisms of Grothendieck groups (induced by pushforward for the first one and pullback for the second one)
$K_{0}(\mathcal{H}^{\mathrm{mon},\mathrm{unip},\mathrm{mix}})=K_{0}(\mathcal{O}^{\mathrm{DL},\mathrm{unip},\mathrm{mix}})=K_{0}(\mathcal{H}^{\mathrm{eq},\mathrm{unip},\mathrm{mix}}).$
Since the pushforward $\mathcal{H}^{\mathrm{mon},\mathrm{unip},\mathrm{mix}}\to\mathcal{O}^{\mathrm{DL},\mathrm{unip},\mathrm{mix}}$ sends standard (resp. costandard, resp. tilting) to standard (resp. costandard, resp. tilting), we deduce from the categorification theorem of loc. cit. that $\nabla_{w,1}^{\mathrm{mix}}]=H_{w^{-1}}^{-1}$ and $T_{w,1}^{\mathrm{mix}}]=\underline{H}_{w}$. Similarly, the pullback $\mathcal{H}^{\mathrm{eq},\mathrm{unip},\mathrm{mix}}\to\mathcal{O}^{\mathrm{DL},\mathrm{unip},\mathrm{mix}}$ sends standard (resp. costandard, resp. simple) to standard (resp. costandard, resp. simple) which allows us to deduce that $\mathrm{IC}_{w,1}^{\mathrm{mix}}]=\underline{\tilde{H}}_{w}$. ∎
Remark 2.5.6.
The reader should note that while there is a ring structure on $\mathcal{H}_{q}$, there is a priori no ring structure or action defined on the RHS. However this equivalence is linear over $\mathbb{Z}[v^{\pm 1}]$.
More generally let $\mathcal{H}_{\mathrm{mon}}$ be the monodromic Hecke algebra defined in [Lus19, 1.4] and see [LY20, Section 3.14] for an account. This is the $\mathbb{Z}[v^{\pm 1}]$-algebra with generators $(T_{w},1_{\chi})$ where $\chi\in\mathrm{Ch}(T)$ subject to the following relations
  1. $(i)$.
    $T_{w}T_{w^{\prime}}=T_{ww^{\prime}}$ if $\ell(w)+\ell(w^{\prime})=\ell(ww^{\prime})$,
  2. $(ii)$.
    $1_{\chi}1_{\chi^{\prime}}=\delta_{\chi,\chi^{\prime}}1_{\chi}$,
  3. $(iii)$.
    $T_{w}1_{\chi}=1_{w\chi}T_{w}$,
  4. $(iv)$.
    $T_{s}^{2}=v^{2}T_{s}+(v^{2}-1)\sum_{s\in\mathbf{W}_{s}^{\circ}}T_{s}1_{\chi}$, we refer to [LY20] for the notation $\mathbf{W}_{s}^{\circ}$.
  5. $(v)$.
    $\sum_{\chi}1_{\chi}=1$, note that combined with the second relation the elements $1_{\chi}$ form a set of orthogonal idempotents and the sum makes sense.
We set $H_{w}=v^{-\ell(w)}T_{w}$ and we note that $\mathcal{H}$ is the direct summand attached to the idempotent $1_{\mathrm{triv}}$. We will also denote by $H_{w,\chi}=H_{w}1_{\chi}$.
As in the unipotent case, this Hecke algebra is equipped with an involution
$\overline{(-)}:\mathcal{H}_{\mathrm{mon}}\to\mathcal{H}_{\mathrm{mon}},v^{n}H_{w}1_{\chi}\mapsto v^{-n}H^{-1}_{w^{-1}}1_{\chi}.$
We refer to [LY20, Section 3.14] for a discussion on the proof of the next theorem.
Theorem 2.5.7.
There exist two self-dual bases $\underline{H}_{w,\chi}$ and $\underline{\tilde{H}}_{w,\chi}$ of $\mathcal{H}_{\mathrm{mon}}$ such that
  1. $(i)$.
    $\underline{H}_{w,\chi}\in H_{w,\chi}+\sum_{y<w,\chi^{\prime}}v\mathbb{Z}[v]H_{y,\chi^{\prime}}$,
  2. $(ii)$.
    $\underline{\tilde{H}}_{w,\chi}\in H_{w,\chi}+\sum_{y<w,\chi^{\prime}}v^{-1}\mathbb{Z}[v^{-1}]H_{y,\chi^{\prime}}$.
As in the monodromic case we have a categorification theorem
Theorem 2.5.8.
There exists an injective map
$K_{0}(\mathcal{O}^{\mathrm{DL},\mathrm{mix}})\to\mathcal{H}_{\mathrm{mon}}$
such that
$\Delta_{w,\chi}^{\mathrm{mix}}]=H_{w,\chi},[\nabla_{w,\chi}^{\mathrm{mix}}]=H_{w^{-1},\chi}^{-1},[\mathrm{IC}_{w,\chi}^{\mathrm{mix}}]=\underline{\tilde{H}}_{w,\chi}.$
Proof.
As in the unipotent case, these statements can be deduced from the corresponding statement from the equivariant categories of [LY20], see in particular Section 3.14 of loc.cit.
Conjecture 2.5.9.
Using the same notation as in Theorem 2.5.8 we have
$T_{w,\chi}^{\mathrm{mix}}]=\underline{H}_{w,\chi}.$
The proof of this conjecture in the unipotent case relied on the study of free monodromic mixed unipotent sheaves as done in [BY13]. We did not find a source in the literature that does this in the non-unipotent mixed case. In the thesis of Gouttard [Gou22], the author studies the non-unipotent case but not in the mixed setting, it can be deduced from the categorification theorems of loc. cit. that
$T_{w,\chi}](1)=\underline{H}_{w,\chi}(1).$
We expect that the generalization of [BY13] to the non-unipotent setting should be a matter of adapting the arguments in the correct framework.
2.6. Relation with category $\mathcal{O}$
The goal of this section is to relate the decomposition numbers appearing in Lemma 2.4.6 to decomposition numbers in the category $\mathcal{O}$. Recall that category $\mathcal{O}$ is the category of sheaves on the stack
$\mathbf{U}\backslash\mathbf{G}/\mathbf{B}.$
For a coefficient ring $\Lambda$, we will denote by
$\mathcal{O}_{\Lambda}=\mathrm{D}(\mathbf{U}\backslash\mathbf{G}/\mathbf{B},\Lambda),$
as in the case of $\mathcal{O}^{\mathrm{DL}}$, we will drop the index $\Lambda$ when the coefficients are clear. We also equip this category with its Bruhat stratification $\mathbf{U}\backslash\mathbf{G}/\mathbf{B}=\sqcup_{w}\mathbf{U}\backslash\mathbf{B}w\mathbf{B}/\mathbf{B}$ and we denote by $j_{w}:\mathbf{U}\backslash\mathbf{B}w\mathbf{B}/\mathbf{B}\subset\mathbf{U}\backslash\mathbf{G}/\mathbf{B}$ the inclusion of the stratum corresponding to $w$. We recall from [AR16], [RSW14] that we have the following objects
  1. $(i)$.
    the standard sheaves $\Delta_{w}^{\mathcal{O},\Lambda}=j_{w,!}\Lambda[\ell(w)]$, where $\Lambda$ denotes the constant sheaf,
  2. $(ii)$.
    the costandard sheaves $\nabla_{w}^{\mathcal{O},\Lambda}=j_{w,*}\Lambda[\ell(w)]$,
  3. $(iii)$.
    the intersection complexes $\mathrm{IC}_{w}^{\mathcal{O},\Lambda}=j_{w,!*}\Lambda[\ell(w)]$,
  4. $(iv)$.
    the indecomposable tilting sheaves $T_{w}^{\mathcal{O},\Lambda}$, which exist by [AR16, Appendix B].
As explained in [AR16, Section 2.7], if $T\in\mathcal{O}^{\heartsuit}$ is a tilting sheaf, then $T[\frac{1}{\ell}]$ is also a tilting sheaf. Consequently, in loc. cit. they introduce multiplicities
$T_{w}^{\mathcal{O},\overline{\mathbb{Z}}_{\ell}}[\frac{1}{\ell}]=T_{w}^{\mathcal{O},\overline{\mathbb{Q}}_{\ell}}\oplus\bigoplus_{v<w}(T_{v}^{\mathcal{O},\overline{\mathbb{Q}}_{\ell}})^{\oplus n^{\mathcal{O}}_{v,w}}.$
Lemma 2.6.1.
For all $v,w\in W$, there is an equality of multiplicities
$n_{v,w}^{\mathcal{O}}=n_{v,w,1}$
where the integer in the RHS is the multiplicity defined in Lemma 2.4.6.
We consider the categories of unipotent free monodromic sheaves on $\mathbf{U}\backslash\mathbf{G}/\mathbf{U}$ as constructed in [BY13] and [Gou22] for $\Lambda=\overline{\mathbb{Q}}_{\ell}$ and $\Lambda=\overline{\mathbb{Z}}_{\ell}$, we denote these categories by $\mathcal{H}^{\mathrm{mon},\mathrm{unip}}_{\overline{\mathbb{Z}}_{\ell}}$ and $\mathcal{H}^{\mathrm{mon},\mathrm{unip}}_{\overline{\mathbb{Q}}_{\ell}}$. Moreover, there is an inversion of $\ell$-functor
$\mathrm{inv}_{\ell}:\mathcal{H}^{\mathrm{mon},\mathrm{unip}}_{\overline{\mathbb{Z}}_{\ell}}\to\mathcal{H}^{\mathrm{mon},\mathrm{unip}}_{\overline{\mathbb{Q}}_{\ell}}$
which is constructed in [Ete24, Section 2.2] essentially as composition of $\overline{\mathbb{Q}}_{\ell}\otimes$$_{\overline{\mathbb{Z}}_{\ell}}$- and a certain completion. We denote by
$\mathfrak{p}^{\prime}:\mathbf{U}\backslash\mathbf{G}/\mathbf{U}\to\mathbf{U}\backslash\mathbf{G}/\mathbf{B}$
the quotient map.
Proof.
Consider the following diagram of categories
$\mathcal{H}^{\mathrm{mon},\mathrm{unip},\overline{\mathbb{Z}}_{\ell}}$$\mathcal{O}^{\mathrm{DL},\overline{\mathbb{Z}}_{\ell}}$$\mathcal{H}^{\mathrm{mon},\mathrm{unip},\overline{\mathbb{Q}}_{\ell}}$$\mathcal{O}^{\overline{\mathbb{Z}}_{\ell}}$$\mathcal{O}^{\mathrm{DL},\overline{\mathbb{Q}}_{\ell}}$$\mathcal{O}^{\overline{\mathbb{Q}}_{\ell}}.$$\scriptstyle{\mathfrak{p}_{!}[\dim\mathbf{T}]}$$\scriptstyle{\mathrm{inv}_{\ell}}$$\scriptstyle{\mathfrak{p}^{\prime}_{!}[\dim\mathbf{T}]}$$\scriptstyle$$\overline{\mathbb{Q}}_{\ell}\otimes$$_{\overline{\mathbb{Z}}_{\ell}}$-$\scriptstyle{\mathfrak{p}_{!}[\dim\mathbf{T}]}$$\scriptstyle{\mathfrak{p}^{\prime}_{!}[\dim\mathbf{T}]}$$\scriptstyle$$\overline{\mathbb{Q}}_{\ell}\otimes$$_{\overline{\mathbb{Z}}_{\ell}}$-}
Let $T\in\mathcal{H}^{\mathrm{mon},\mathrm{unip},\overline{\mathbb{Z}}_{\ell}}$ be a free monodromic tilting object, then the object $\mathrm{inv}_{\ell}T\in\mathcal{H}^{\mathrm{mon},\mathrm{unip},\overline{\mathbb{Q}}_{\ell}}$ also free monodromic tilting. Furthermore the following relations hold by definition of the functor $\mathrm{inv}_{\ell}$. We have
$\mathfrak{p}_{!}\mathrm{inv}_{\ell}T[\dim\mathbf{T}]=((\mathfrak{p}_{!}T)[\frac{1}{\ell}])^{s=1}[\dim\mathbf{T}]$
where unip denotes the direct summand of $(p_{!}T)[\frac{1}{\ell}])$ corresponding to $s=1$ in the decomposition of lemma 2.4.6. We also have
$\mathfrak{p}^{\prime}_{!}\mathrm{inv}_{\ell}T[\dim\mathbf{T}]=(\mathfrak{p}^{\prime}_{!}T)[\frac{1}{\ell}][\dim\mathbf{T}].$
Let $\hat{T}_{w}^{\overline{\mathbb{Z}}_{\ell}}$ be the indcomposable free monodromic tilting sheaf in $\mathcal{H}^{\mathrm{mon},\mathrm{unip},\overline{\mathbb{Z}}_{\ell}}$ corresponding to $w$. By Remark 2.3.4, we have
$\mathfrak{p}_{!}\hat{T}_{w}^{\overline{\mathbb{Z}}_{\ell}}[\dim\mathbf{T}]=T_{w,1}^{\overline{\mathbb{Z}}_{\ell}}$
and by [BR22, Lemma 5.8], we have
$\mathfrak{p}^{\prime}_{!}\hat{T}_{w}^{\overline{\mathbb{Z}}_{\ell}}[\dim\mathbf{T}]=T_{w}^{\mathcal{O},\overline{\mathbb{Z}}_{\ell}}.$
Now since the category of free monodromic tilting sheaves in $\mathcal{H}^{\mathrm{mon},\mathrm{unip},\overline{\mathbb{Q}}_{\ell}}$ is Krull-Schmidt there is a decomposition
$\mathrm{inv}_{\ell}\hat{T}_{w}^{\overline{\mathbb{Z}}_{\ell}}=\hat{T}^{\overline{\mathbb{Q}}_{\ell}}_{w}\oplus\bigoplus_{v<w}(\hat{T}^{\overline{\mathbb{Q}}_{\ell}}_{v})^{\oplus\hat{n}_{v,w}}.$
It now follows that $n_{v,w,1}=\hat{n}_{v,w}=n^{\mathcal{O}}_{v,w}$. ∎
Remark 2.6.2.
The decomposition numbers that appear in Lemma 2.6.1 control the difference between the usual Kazhdan–Lusztig basis in $\mathcal{H}$ and the $\ell$-Kazhdan–Lusztig basis (which depends on $\mathbf{G}$ and not just on $\mathbf{W}$), firstly defined in the case of crystallographic Coxeter groups in [JMW14]. We refer to [JW17] for an account on this basis.
Lemma 2.6.3.
For a fixed $\mathbf{G}$, there exists $\ell_{0}$ large enough such that for all $\ell\geq\ell_{0}$ the Kazhdan-Lusztig and $\ell$-Kazhdan–Lusztig basis coincide. For such $\ell$, we have $n_{v,w,1}=0$ for all $v<w$.
Proof.
By [AR16, Theorem 2.6], if $\ell$ is good for $\mathbf{G}$, these decomposition numbers are equal to decomposition numbers of parity sheaves. By [JMW14, Proposition 2.41], mod $\ell$ indecomposable parity sheaves and $\mathrm{IC}$ sheaves coincide for all but finitely many primes $\ell$. ∎
Remark 2.6.4.
Predicting the behaviour of this basis and in particular when it differs from the usual Kazhdan–Lusztig basis is a very hard problem. In [GJW23], the authors construct an algorithm to compute this basis.
3. Tilting representations
3.1. Horocycle correspondence
We introduce the horocycle correspondence which we attribute to Lusztig. This same correspondence (not presented in a stacky and twisted way) was used to first define character sheaves [Lus85].
$\frac{\mathbf{G}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{G}}\xleftarrow{\alpha}\frac{\mathbf{G}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{B}}\xrightarrow{\beta}\frac{\mathbf{U}\backslash\mathbf{G}/\mathbf{U}}{\mathrm{Ad}_{\mathrm{F}}\mathbf{T}}.$
We introduce the following functors (which stand for horocycle and character)
$\mathrm{ch}=\alpha_{!}\beta^{*}[\dim\mathbf{U}](\frac{\dim\mathbf{U}}{2}):\mathcal{O}^{\mathrm{DL}}\to\mathrm{D}(\mathrm{Rep}_{\Lambda}\mathbf{G}^{\mathrm{F}})$
and
$\mathrm{hc}=\alpha_{!}\beta^{*}[-\dim\mathbf{U}](-\frac{\dim\mathbf{U}}{2}):\mathrm{D}(\mathrm{Rep}_{\Lambda}\mathbf{G}^{\mathrm{F}})\to\mathcal{O}^{\mathrm{DL}}.$
Note that since $\alpha$ is proper and smooth, and $\beta$ is smooth the functor $\mathrm{ch}$ is right adjoint to $\mathrm{hc}$.
Theorem 3.1.1 ([Ete24, Corollary 3.3.5, Lemma 3.2.7]).
  1. $(i)$.
    The functor $\mathrm{hc}$ is conservative.
  2. $(ii)$.
    There is an isomorphism of functors $\mathrm{ch}j_{w,!}=R_{w}[\ell(w)]$ where the RHS is the Deligne–Lusztig induction functor.
Let us also record the compatibility with (geometric) Lusztig series.
Theorem 3.1.2.
  1. $(i)$.
    The category $\mathrm{D}(\mathrm{Rep}_{\Lambda}\mathbf{G}^{\mathrm{F}})$ splits as
    $\mathrm{D}(\mathrm{Rep}_{\Lambda}\mathbf{G}^{\mathrm{F}})=\bigoplus_{s\in(G_{\Lambda}^{*}/\mathrm{ss})^{\mathrm{F}^{*}}}\mathrm{D}(\mathrm{Rep}^{s}_{\Lambda}\mathbf{G}^{\mathrm{F}})$
    where $\mathrm{D}(\mathrm{Rep}^{s}_{\Lambda}\mathbf{G}^{\mathrm{F}})$ is the Lusztig series corresponding to the conjugacy class $s$.
  2. $(ii)$.
    The functors $\mathrm{ch}$ and $\mathrm{hc}$ preserve the splittings into Lusztig series and the splitting of $\mathcal{O}^{\mathrm{DL}}$ of Corollary 2.4.4.
Proof.
The first point is done in the original paper of Deligne and Lusztig [DL76] if $\Lambda=\overline{\mathbb{Q}}_{\ell}$. The integral and the modular cases are done in [BM89, Theorem 2.2]. In view of Theorem 3.1.1, the second point follows from the definition of the Lusztig series and the compatibility between the Deligne–Lusztig functors and the functors $\mathrm{ch}$ and $\mathrm{hc}$ of theorem 3.1.1. ∎
3.2. Tilting representations
Definition 3.2.1.
A tilting representation of $\mathbf{G}^{\mathrm{F}}$ is an object in $\mathrm{D}(\mathrm{pt}/\mathbf{G}^{\mathrm{F}},\Lambda)$ of the form $\mathrm{ch}(T)$ where $T\in\mathcal{O}_{\mathrm{tilt}}^{\mathrm{DL}}$.
Lemma 3.2.2.
All the complexes $\mathrm{ch}(\bigoplus_{\chi}\Delta_{w,\chi})$ are concentrated in nonnegative degree.
Proof.
This is a consequence of Artin vanishing as all Deligne–Lusztig varieties are affine. If $q$ is large enough then this is [DL76, Theorem 9.7], the general case is done in [BHL25, Corollary 3.12]. ∎
Remark 3.2.3.
Another proof of Lemma 3.2.2 is done in [Zhu25] and is independent of [BHL25].
Theorem 3.2.4.
 [^1] This result was obtained independently by Zhu [Zhu25]
Let $T\in\mathcal{O}_{\mathrm{tilt}}^{\mathrm{DL}}$.
  1. $(i)$.
    The complex $\mathrm{ch}(T)$ is a compact object of $\mathrm{D}(\mathrm{Rep}_{\Lambda}\mathbf{G}^{\mathrm{F}})$.
  2. $(ii)$.
    The complex $\mathrm{ch}(T)$ is concentrated in degree 0 and is a projective object of $\mathrm{Rep}_{\Lambda}\mathbf{G}^{\mathrm{F}}$.
Proof.
Since the map $\alpha$ is proper and smooth and the map $\beta$ is smooth the functor $\mathrm{ch}$ has a continuous right adjoint hence $\mathrm{ch}$ preserves compact objects. Since tilting sheaves are compact by Proposition 2.1.3, the objects $\mathrm{ch}(T)$ are compact.
By Theorem 3.1.1 and Lemma 3.2.2, it follows that $\mathrm{ch}(\Delta)\in\mathrm{D}^{\geq 0}(\mathrm{pt}/\mathbf{G}^{\mathrm{F}})$ (resp. $\mathrm{ch}(\nabla)\in\mathrm{D}^{\leq 0}(\mathrm{pt}/\mathbf{G}^{\mathrm{F}})$). Since $T$ has both a $\Delta$ and a $\nabla$-filtration, we get that $\mathrm{ch}(T)\in\mathrm{D}^{\heartsuit}(\mathrm{pt}/\mathbf{G}^{\mathrm{F}})$. Hence $\mathrm{ch}(T)$ is a representation of $\mathbf{G}^{\mathrm{F}}$ concentrated in degree 0, which is compact in $\mathrm{D}(\mathrm{pt}/\mathbf{G}^{\mathrm{F}})$. Since $\Lambda[\mathbf{G}^{\mathrm{F}}]$ is a symmetric algebra, it follows that the representation $\mathrm{ch}(T)$ is projective. ∎
Proposition 3.2.5.
The representations $\mathrm{ch}(T)$ for $T\in\mathcal{O}_{\mathrm{tilt}}^{\mathrm{DL}}$ generate the triangulated category $\mathrm{Perf}(\Lambda[\mathbf{G}^{\mathrm{F}}])$.
Proof.
Since the $\mathrm{ch}(T)$ generate the category $\mathcal{O}_{\mathrm{tilt}}^{\mathrm{DL}}$ this statement is equivalent to the conservativity property of the functor $\mathrm{hc}$ of Theorem 3.1.1. ∎
4. Application to Alvis–Curtis duality
4.1. The Alvis–Curtis dual of an $\mathrm{IC}$ representation
We consider the category $\mathrm{D}_{\mathrm{c}}(\mathrm{pt}/\mathbf{G}^{\mathrm{F}},\overline{\mathbb{Q}}_{\ell})^{\mathrm{mix}}$ of representations of $\mathbf{G}^{\mathrm{F}}$ such that the underlying complex of vector spaces is mixed, where we use, as in Section 2.5, the Frobenius $\mathrm{F}^{\delta}$.
We consider the Alvis–Curtis duality functor, see [DL82]
$d:\mathrm{D}_{\mathrm{c}}(\mathrm{pt}/\mathbf{G}^{\mathrm{F}},\overline{\mathbb{Q}}_{\ell})^{\mathrm{mix}}\to\mathrm{D}_{\mathrm{c}}(\mathrm{pt}/\mathbf{G}^{\mathrm{F}},\overline{\mathbb{Q}}_{\ell})^{\mathrm{mix}}.$
We denote by $R(\mathbf{G}^{\mathrm{F}})$ the Grothendieck group of $\mathrm{D}_{\mathrm{c}}(\mathbf{G}^{\mathrm{F}},\overline{\mathbb{Q}}_{\ell})$ and we consider the ring $\mathbb{Z}[v^{\pm 1}]\otimes R(\mathbf{G}^{\mathrm{F}})$. There is natural map
$\gamma:K_{0}(\mathrm{D}_{\mathrm{c}}(\mathrm{pt}/\mathbf{G}^{\mathrm{F}},\overline{\mathbb{Q}}_{\ell})^{\mathrm{mix}})\to\mathbb{Z}[v^{\pm 1}]\otimes R(\mathbf{G}^{\mathrm{F}})$
which sends the class of a representation $\rho=\oplus_{j}\rho_{j}$ where $\rho_{j}$ is the generalized eigenspace of Frobenius of weight $j$ to $\sum_{j}v^{j}\otimes[\rho_{j}]$. The group $K_{0}(\mathrm{D}_{\mathrm{c}}(\mathrm{pt}/\mathbf{G}^{\mathrm{F}},\overline{\mathbb{Q}}_{\ell})^{\mathrm{mix}})$ is $\mathbb{Z}[v^{\pm 1}]$-linear where $v^{-1}$ acts by the half-Tate twist and the map $\gamma$ is $\mathbb{Z}[v^{\pm 1}]$-linear.
Remark 4.1.1.
The functor $\gamma\circ\mathrm{ch}$ induces a $\mathbb{Z}[v^{\pm 1}]$-linear morphism
$K_{0}(\mathcal{O}^{\mathrm{DL},\mathrm{mix}}_{\overline{\mathbb{Q}}_{\ell}})\to\mathbb{Z}[v^{\pm 1}]\otimes R(\mathbf{G}^{\mathrm{F}}).$
Let $a:\mathcal{H}\to\mathcal{H}$ be the automorphism of the Hecke algebra given by
$a(v)=v,a(H_{x})=(-1)^{\ell(x)}H_{x^{-1}}^{-1}.$
Note that we have $\underline{H}_{w}=(-1)^{\ell(w)}a(\underline{\tilde{H}}_{w})$, see [Soe97, Theorem 2.7].
We will denote, as in [Lus84, Section 3.1], $\mathrm{Irr}(\mathbf{W})_{\mathrm{ex}}$ the set of irreducible modules of $\mathbf{W}$ which can be extended to a $\mathbf{W}\rtimes\langle\mathrm{F}\rangle$-module, and for $\widetilde{E}\in\mathrm{Irr}(\mathbf{W})_{\mathrm{ex}}$ we denote by $\widetilde{E}(v)$ the corresponding $\mathbb{Q}(v)$-irreducible $\mathcal{H}\otimes_{\mathbb{Z}[v^{\pm 1}]}\mathbb{Q}(v)$-representation. Finally given $\widetilde{E}\in\mathrm{Irr}(\mathbf{W})_{\mathrm{ex}}$, we denote by $R_{\widetilde{E}}$ the corresponding almost character of $\mathbf{G}^{\mathrm{F}}$ constructed in [Lus84] and we consider the following two maps $\mathcal{H}\to\mathbb{Q}(v)\otimes R(\mathbf{G}^{\mathrm{F}})$
  1. $(i)$.
    $\mathcal{H}=K_{0}(\mathcal{O}^{\mathrm{DL},\mathrm{mix},\mathrm{unip}}_{\overline{\mathbb{Q}}_{\ell}})\xrightarrow{\mathrm{ch}}K_{0}(\mathrm{Rep}^{\mathrm{mix}}_{\overline{\mathbb{Q}}_{\ell}}\mathbf{G}^{\mathrm{F}})\xrightarrow{\gamma}\mathbb{Q}(v)\otimes R(\mathbf{G}^{\mathrm{F}})$, we shall denote this map by $\mathrm{ch}$,
  2. $(ii)$.
    $\mathrm{tr}:\mathcal{H}\to\mathbb{Q}(v)\otimes R(\mathbf{G}^{\mathrm{F}})$ the map defined by $\mathrm{tr}(f)=\sum_{\widetilde{E}}\mathrm{tr}(f\mathrm{F},\widetilde{E}(u))\otimes R_{\tilde{E}}$.
Theorem 4.1.2 ([Lus84, Theorem 3.8]).
There is an equality of $\mathbb{Z}[v^{\pm 1}]$-linear maps $\mathcal{H}\to\mathbb{Q}(v)\otimes R(\mathbf{G}^{\mathrm{F}})$
$\mathrm{tr}=\mathrm{ch}\circ b.$
(Recall that $b$ is defined in Remark 2.5.4.)
Proof.
Tracing the definitions, the statement of [Lus84, Theorem 3.8] is precisely this equality evaluated at $\underline{H}_{w}$. Since these elements form a basis of $\mathcal{H}$, the theorem follows. ∎
Theorem 4.1.3 ([Lus84, Proposition 6.9]).
There is an equality in $\mathbb{Q}(v)\otimes R(\mathbf{G}^{\mathrm{F}})$,
$d\mathrm{ch}(\underline{\tilde{H}}_{w,1})=\pm\mathrm{tr}(\underline{\tilde{H}}_{w,1}).$
The signs are determined in loc. cit., we will not need them so we do not introduce them to avoid cumbersome notations.
Using the categorification Theorem 2.5.5, we then deduce the following corollary.
Corollary 4.1.4.
There is an equality in $\mathbb{Q}(v)\otimes R(\mathbf{G}^{\mathrm{F}})$
$d(\mathrm{ch}([\mathrm{IC}_{w,1}^{\mathrm{mix}}]))=\pm\mathrm{ch}([T_{w,1}^{\mathrm{mix}}]).$
Remark 4.1.5.
We have restricted ourselves to the unipotent case, as this suffices to prove the conjecture of Dudas–Malle. However it should be noted that, assuming Conjecture 2.5.9, a non-unipotent variant of Corollary 4.1.4 could be deduced. Indeed, Theorems 4.1.2 and 4.1.3 are both shown in [Lus84] in the non-unipotent case.
4.2. Eigenvalues of Frobenius
We make a quick reminder on the theory of eigenvalues of Frobenius.
Theorem 4.2.1 ([Lus84, Theorem 3.8.1]).
Let $\rho\in\mathrm{Irr}_{\overline{\mathbb{Q}}_{\ell}}^{\mathrm{unip}}\mathbf{G}^{\mathrm{F}}$ be an irreducible unipotent representation. There exists a root of unity $\zeta_{\rho}\in\overline{\mathbb{Q}}_{\ell}^{\times}$ satisfying the following property, for all $A\in\mathcal{O}^{\mathrm{DL},\mathrm{unip},\mathrm{mix}}$, the eigenvalues of $\mathrm{F}^{\delta}$ on the $\rho$-isotypic component of $\mathrm{ch}(A)$, are of the form $q^{\frac{i}{2}}\zeta$ for $i\in\mathbb{Z}$.
Proof.
This exact statement is proved in [Lus84, Theorem 3.8.1] when $A=\Delta_{w,1}^{\mathrm{mix}}$ (and thus $\mathrm{ch}(A)=\mathrm{R\Gamma}_{c}(X(w),\overline{\mathbb{Q}}_{\ell})[\ell(w)]$). Since these objects generate the category $\mathcal{O}^{\mathrm{DL},\mathrm{unip},\mathrm{mix}}$, the statement follows. ∎
Let $\rho$ be as in Theorem 4.2.1, and let $d(\rho)\in\mathrm{Irr}_{\overline{\mathbb{Q}}_{\ell}}^{\mathrm{unip}}\mathbf{G}^{\mathrm{F}}$ be the Alvis–Curtis dual of $\rho$ (written without signs).
Lemma 4.2.2.
The following equality holds
$\zeta_{d(\rho)}=\zeta_{\rho}.$
Proof.
By [Lus78, 3.33] (see also [GM03, Lemma 4.2]), the map
$\rho\mapsto\zeta_{\rho}$
is constant on Harish-Chandra series. Since Harish-Chandra series are invariant under Alvis–Curtis duality, the statement follows. ∎
Recall (see 1.3) that we denote by $\mathrm{D}_{c}(\mathrm{pt}/\mathbf{G}^{\mathrm{F}},\overline{\mathbb{Q}}_{\ell})^{\mathrm{Weil}}$ the category of pairs $(\rho,\alpha)$ where $\rho\in\mathrm{D}_{c}(\mathrm{pt}/\mathbf{G}^{\mathrm{F}},\overline{\mathbb{Q}}_{\ell})$ and $\alpha$ is an isomorphism $\mathrm{F}^{\delta,*}\rho\simeq\rho$. For $\lambda\in\overline{\mathbb{Q}}_{\ell}$ and $(\rho,\alpha)\in\mathrm{D}_{c}(\mathrm{pt}/\mathbf{G}^{\mathrm{F}},\overline{\mathbb{Q}}_{\ell})^{\mathrm{Weil}}$, we denote by
$\rho[\lambda]$
the direct summand of $(\rho,\alpha)$ where $\alpha$ has generalized eigenvalue $\lambda$. For $\overline{\lambda}\in\overline{\mathbb{F}}_{\ell}$, we denote by
$\rho[\overline{\lambda}]=\bigoplus_{\lambda}\rho[\lambda]$
where the direct sum is indexed by all $\lambda\in\overline{\mathbb{Z}}_{\ell}$ which reduce to $\overline{\lambda}\in\overline{\mathbb{F}}_{\ell}$.
Lemma 4.2.3.
Let $A\in\mathcal{O}^{\mathrm{DL},\mathrm{unip},\mathrm{mix}}$, then there is an equality in $K_{0}(\mathrm{Rep}_{\overline{\mathbb{Q}}_{\ell}}(\mathbf{G}^{\mathrm{F}}))$ for all $\lambda\in\overline{\mathbb{Q}}_{\ell}^{\times}$,
$d(\mathrm{ch}(A)[\lambda])=(d(\mathrm{ch}(A))[\lambda].$
Proof.
It is enough to prove the statement for a generating collection of sheaves in $\mathcal{O}^{\mathrm{DL},\mathrm{unip},\mathrm{mix}}$, we will prove it for $\mathrm{IC}_{w}^{\mathrm{mix}}$. By Theorem 4.2.1, we can assume that $\lambda=q^{\frac{i}{2}}\zeta$ for some root of unity $\zeta$. Since the equality of Corollary 4.1.4 is $\mathbb{Q}(v)$-linear, we have
$\sum_{\zeta^{\prime}}d(\mathrm{IC}_{w}^{\mathrm{mix}}[q^{\frac{i}{2}}\zeta^{\prime}])=\sum_{\zeta^{\prime}}d(\mathrm{IC}_{w}^{\mathrm{mix}})[q^{\frac{i}{2}}\zeta^{\prime}].$ (1)
Let $\rho\in\mathrm{Irr}_{\overline{\mathbb{Q}}_{\ell}}^{\mathrm{unip}}\mathbf{G}^{\mathrm{F}}$, applying $\langle\rho,$-$\rangle$ to 1, we have
$\sum_{\zeta^{\prime}}\langle\rho,d(\mathrm{IC}_{w}^{\mathrm{mix}}[q^{\frac{i}{2}}\zeta^{\prime}])\rangle=\sum_{\zeta^{\prime}}\langle\rho,d(\mathrm{IC}_{w}^{\mathrm{mix}})[q^{\frac{i}{2}}\zeta^{\prime}]\rangle.$ (2)
By Theorem 4.2.1, the LHS is
$\sum_{\zeta^{\prime}}\langle\rho,d(\mathrm{IC}_{w}^{\mathrm{mix}}[q^{\frac{i}{2}}\zeta^{\prime}])\rangle$ $=\sum_{\zeta^{\prime}}\langle d(\rho),\mathrm{IC}_{w}^{\mathrm{mix}}[q^{\frac{i}{2}}\zeta^{\prime}]\rangle$
$=\langle d(\rho),\mathrm{IC}_{w}^{\mathrm{mix}}[q^{\frac{i}{2}}\zeta_{d(\rho)}]\rangle$
$=\langle\rho,d(\mathrm{IC}_{w}^{\mathrm{mix}}[q^{\frac{i}{2}}\zeta_{d(\rho)}])\rangle.$
Similarly, the RHS evaluates to
$\sum_{\zeta^{\prime}}\langle\rho,d(\mathrm{IC}_{w}^{\mathrm{mix}})[q^{\frac{i}{2}}\zeta^{\prime}]\rangle=\langle\rho,d(\mathrm{IC}_{w}^{\mathrm{mix}})[q^{\frac{i}{2}}\zeta_{\rho}]\rangle.$
It thus follows from Lemma 4.2.2 that
$\langle\rho,d(\mathrm{IC}_{w}^{\mathrm{mix}}[q^{\frac{i}{2}}\zeta])\rangle=\langle\rho,d(\mathrm{IC}_{w}^{\mathrm{mix}})[q^{\frac{i}{2}}\zeta]\rangle$
for all $\zeta$ and the lemma follows. ∎
Corollary 4.2.4.
Let $A\in\mathcal{O}^{\mathrm{DL},\mathrm{unip},\mathrm{mix}}$, for all $\overline{\lambda}\in\overline{\mathbb{F}}_{\ell}^{\times}$, there is an equality in $K_{0}(\mathrm{Rep}_{\overline{\mathbb{Q}}_{\ell}}(\mathbf{G}^{\mathrm{F}}))$
$d(\mathrm{ch}(A)[\overline{\lambda}])=(d(\mathrm{ch}(A))[\overline{\lambda}]).$
Proof.
This is simply a matter of grouping terms and applying Lemma 4.2.3. ∎
4.3. A conjecture of Dudas–Malle
The goal of this section is to provide an answer to a conjecture of Dudas and Malle [DM14, Conjecture 1.2]. The following hypothesis is in force in this section :
The Kazhdan–Lusztig basis and the $\ell$-Kazhdan–Lusztig basis coincide.
By Lemma 2.6.3 this is guaranteed whenever $\ell$ is large enough (depending only on $\mathbf{G}$ and not on the Frobenius).
Let $(w,\chi)$ be a pair with $w\in\mathbf{W}$ and $\chi$ a $\overline{\mathbb{Q}}_{\ell}$-character of $\mathbf{T}^{w\mathrm{F}}$. The object $\mathrm{IC}_{w,\chi}^{\mathrm{mix}}$ is equipped with a canonical Weil-structure, i.e., an isomorphism
$\mathrm{F}^{\delta,*}\mathrm{IC}_{w,\chi}^{\mathrm{mix}}\simeq\mathrm{IC}_{w,\chi}^{\mathrm{mix}}.$
By functoriality and compatibility with the Frobenius, the object $\mathrm{ch}(\mathrm{IC}_{w,\chi}^{\mathrm{mix}})$ is also equipped with an isomorphism
$\mathrm{F}^{\delta,*}\mathrm{ch}(\mathrm{IC}_{w,\chi}^{\mathrm{mix}})\simeq\mathrm{ch}(\mathrm{IC}_{w,\chi}^{\mathrm{mix}}).$
Theorem 4.3.1 ([DM14, Conjecture 1.2]).
For all $\overline{\lambda}\in\overline{\mathbb{F}}_{\ell}$, the character
$d(\mathrm{ch}(\mathrm{IC}_{w,1}^{\mathrm{mix}}[\overline{\lambda}]))$
is the unipotent part of a character of a projective $\overline{\mathbb{Z}}_{\ell}$-representation.
Proof.
Consider the object $T_{w,1}^{\overline{\mathbb{Z}}_{\ell}}$. We have
$T_{w,1}^{\overline{\mathbb{Z}}_{\ell}}[\frac{1}{\ell}]=T_{w,1}^{\overline{\mathbb{Q}}_{\ell}}\oplus\bigoplus_{\chi_{\ell}\neq 1}T_{w,\chi_{\ell}}^{\overline{\mathbb{Q}}_{\ell}}\oplus\bigoplus_{v<w,\chi_{\ell}\neq 1}T_{v,\chi_{\ell}}^{\overline{\mathbb{Q}}_{\ell},\oplus{n_{v,w,\chi_{\ell}}}}$
by Lemmas 2.4.6 and 2.6.3. We equip $T_{w,1}^{\overline{\mathbb{Z}}_{\ell}}[\frac{1}{\ell}]$ with a Weil structure such that there is an isomorphism of sheaves with Weil structures
$T_{w,1}^{\overline{\mathbb{Z}}_{\ell}}[\frac{1}{\ell}]=T_{w,1}^{\overline{\mathbb{Q}}_{\ell},\mathrm{mix}}\oplus\bigoplus_{\chi_{\ell}\neq 1}T_{w,\chi_{\ell}}^{\overline{\mathbb{Q}}_{\ell},\mathrm{mix}}\oplus\bigoplus_{v<w,\chi_{\ell}\neq 1}T_{v,\chi_{\ell}}^{\overline{\mathbb{Q}}_{\ell},\mathrm{mix},\oplus{n_{v,w,\chi_{\ell}}}}.$
Applying $\mathrm{ch}(-)[\overline{\lambda}]$ we get an isomorphism
$\mathrm{ch}(T_{w,1}^{\overline{\mathbb{Z}}_{\ell}}[\frac{1}{\ell}])[\overline{\lambda}]=\mathrm{ch}(T_{w,1}^{\overline{\mathbb{Q}}_{\ell},\mathrm{mix}})[\overline{\lambda}]\oplus\bigoplus_{\chi_{\ell}\neq 1}\mathrm{ch}(T_{w,\chi_{\ell}}^{\overline{\mathbb{Q}}_{\ell},\mathrm{mix}})[\overline{\lambda}]\oplus\bigoplus_{v<w,\chi_{\ell}\neq 1}\mathrm{ch}(T_{v,\chi_{\ell}}^{\overline{\mathbb{Q}}_{\ell},\mathrm{mix},\oplus{n_{v,w,\chi_{\ell}}}})[\overline{\lambda}].$
The unipotent part of this representation is $\mathrm{ch}(T_{w,1}^{\overline{\mathbb{Q}}_{\ell},\mathrm{mix}})[\overline{\lambda}]$. Since mixed tilting sheaves are filtered, as sheaves with Weil structure, by mixed standard sheaves, it follows that the only eigenvalues of Frobenius appearing in $\mathrm{ch}(T_{w,\chi}^{\overline{\mathbb{Q}}_{\ell},\mathrm{mix}})$ are eigenvalues of Frobenius appearing on the cohomology of Deligne–Lusztig varieties. By Theorem 4.2.1, these eigenvalues of Frobenius are of the form $q^{\frac{i}{2}}\zeta$ where $i\in\mathbb{Z}$ and $\zeta\in\overline{\mathbb{Q}}_{\ell}^{\times}$ are roots of unity. It follows that the Weil structure on $\mathrm{ch}(T_{w,1}^{\overline{\mathbb{Z}}_{\ell}}[\frac{1}{\ell}])$, which is an isomorphism
$\mathrm{F}^{\delta,*}\mathrm{ch}(T_{w,1}^{\overline{\mathbb{Z}}_{\ell}}[\frac{1}{\ell}])\simeq\mathrm{ch}(T_{w,1}^{\overline{\mathbb{Z}}_{\ell}}[\frac{1}{\ell}])$
descends to $\overline{\mathbb{Z}}_{\ell}$ as all its eigenvalues lie in $\overline{\mathbb{Z}}_{\ell}$. Hence we have a well-defined $\overline{\mathbb{Z}}_{\ell}$-linear isomorphism
$\mathrm{F}^{\delta,*}\mathrm{ch}(T_{w,1}^{\overline{\mathbb{Z}}_{\ell}})\simeq\mathrm{ch}(T_{w,1}^{\overline{\mathbb{Z}}_{\ell}}).$
There is therefore a well-defined direct summand of $\mathrm{ch}(T_{w,1}^{\overline{\mathbb{Z}}_{\ell}})$ which we denote by $\mathrm{ch}(T_{w,1}^{\overline{\mathbb{Z}}_{\ell}})[\overline{\lambda}]$ over $\overline{\mathbb{Z}}_{\ell}$ and which is characterized by the property that
$\mathrm{ch}(T_{w,1}^{\overline{\mathbb{Z}}_{\ell}})[\overline{\lambda}][\frac{1}{\ell}]=\mathrm{ch}(T_{w,1}^{\overline{\mathbb{Z}}_{\ell}})[\frac{1}{\ell}][\overline{\lambda}].$
Since $\mathrm{ch}(T_{w,1}^{\overline{\mathbb{Z}}_{\ell}})[\overline{\lambda}]$ is a direct summand of $\mathrm{ch}(T_{w,1}^{\overline{\mathbb{Z}}_{\ell}})$, by Theorem 3.2.4, this is a projective representation of $\mathbf{G}^{\mathrm{F}}$ over $\overline{\mathbb{Z}}_{\ell}$. To prove the theorem, it is enough to show that, after taking characters, we have
$d(\mathrm{ch}(T_{w,1}^{\overline{\mathbb{Q}}_{\ell},\mathrm{mix}}[\overline{\lambda}]))=\mathrm{ch}(\mathrm{IC}_{w,1}^{\overline{\mathbb{Q}}_{\ell},\mathrm{mix}})[\overline{\lambda}].$
This is exactly Corollary 4.2.4. ∎